### v2

Abstract
In this paper I develop and characterize a random attention model with unobserved reference alternatives. The decision maker pays attention to different subsets of the available set of alternatives randomly. The reference alternatives are exactly those alternatives that are always paid attention to, i.e. they are attention-privileged and these alternatives are supposed to be unknown to the outside observer. The characterization allows for a complete identification of the reference alternatives and a coarse identification of the underlying preferences. I then restrict the model by considering the independent random attention function and provide a complete identification of the underlying preferences as well.
JEL Classification: D01, D91, Z13
Keywords: Bounded Rationality, Stochastic Choice, Limited Attention, Consideration Sets
1 Introduction
Decision-making is often constrained by limited attention, whether due to limits to cognitive processing because of too many alternatives to choose from or due to lack of time. In such situations, decision-makers (DM) often resort to heuristics that can help them decide relatively quickly with less mental exercise. If the DM has to make similar choices multiple times under the same primitives, the heuristic that decides how to allocate attention can lead to different choices. Stochastic choice functions represent such choices with the main research task then been to ascertain the likely underlying heuristic and the primitives used by the DM. Cattaneo et al. (2020) introduced the framework of random attention models (RAM) to deal with the analysis of stochastic choice functions with random choice due to limited attention.
There has been plenty of literature dealing with observable status quo affecting the choices of a DM. In particular, Kovach and Suleymanov (2023) uses the framework of RAM to look at how a reference alternative, or a status quo in their case, can affect choices if it directs attention. These authors were the first to deal with observable reference alternatives’ impact on attention in a stochastic choice setting. It is natural to presume that a reference alternative can direct attention when the underlying heuristic is applied due to limited attention, for which they provide a characterization of a reference-dependent RAM, as well as special cases where the attention function is further restricted.
In certain cases, it may make sense that the reference is observable. For instance, for a decision maker in the market for a new car, his current car can be a reference. However, in many situations, it is not quite realistic to assume the knowledge of the DM’s reference to anyone else but the DM himself. Secondly, to analyse a choice problem with observable references, we have to deal with a family of stochastic choice functions, one for each alternative as the reference, which is quite a large data requirement. If we restrict the references to be unknown we deal with only menus and stochastic choice functions over alternatives and menus. We do not have to deal with families of stochastic choice functions for each possible reference.
It is also not obvious that a reference alternative has to be unique. In what follows we allow for multiple reference alternatives; clearly, one particular such alternative could be the status quo. In this spirit, I define an alternative to be a reference alternative if and only if it is attention privileged. However, since I am looking at choices from menus that depend on reference alternatives, I assume that the references are menu-independent. It is certainly plausible that references can be menu-dependent, through which we can explain a larger set of stochastic choice behavior, but dealing with menu-dependent unobservable references is beyond the scope of this paper. I provide below two examples that match the mentioned assumptions:
  • Example 1: Michael Jordan is associated with the Air Jordan brand, a fan of him who is interested in buying shoes might consider an Air Jordan shoe as a reference.
  • Example 2: Suppose your neighbor buys a new flat-screen 4K TV and invites you over to watch a football game together. You are startled by the picture quality and ask about the product. The prior knowledge of this particular TV will mean that this product will act as a reference point the next time you buy a TV.
In the above examples, it is not obvious that the reference will be observable to the outside observer. Neither is there any evidence that the reference is unique, a person could be a fan of multiple athletes each with their own clothing line, or have multiple suggestions when deciding to buy a new TV.
In this paper, I look at a random attention model with unobservable menu-independent reference alternatives that are attention-privileged (RAM-UR). I provide a characterization of the RAM-UR by utilizing three axioms and identifying the corresponding primitives (i.e., the set of reference alternatives and the underlying preferences) We can uniquely identify the set of references but the underlying preferences are only coarsely identified, i.e. the identification is with respect to the relative ranking between any alternative and any reference alternative and the transitive closure thereof. I then restrict the attention function to be of the independent random attention (IRA) form (Manzini and Mariotti (2014)) and show that under such a restriction, the preferences can be fully identified. The existence of menus with no references is crucial in making the preference identification in this case which is possible here because of the menu-independence of the references. Kovach and Suleymanov (2023) allowed for the references to be part of the primitive which precludes them from considering menus without references. I also elaborate on the fact that rational choice is a boundary case of the RAM-UR as well as the RAM-UR-IRA, and the IRA model is also a boundary case of the RAM-UR-IRA [^3] Boundary cases can be of two types: When the set of references is the set of all alternatives, or when the set of references is empty..
The rest of the paper is organized as follows. In Section 2, I introduce the general RAM-UR model, axiomatically characterize it, and discuss the identification results. The characterization of the RAM-UR-IRA is then provided in Section 3, where I also discuss the nestedness properties of the corresponding random attention models. Section 4 relates my work to the existing literature. All proofs are relegated to Appendix 1.
2 Random Attention Model with Unobserved References
2.1 Primitives
Let $X$ be the finite set of all possible alternatives. Denote the set of all non-empty subsets of $X$ by $\mathcal{X}$, with a typical element being $A$.
Definition 1
A stochastic choice function is a mapping $p:X\times\mathcal{X}\rightarrow[0,1]$ where for each $A\in\mathcal{X}$ the following three conditions hold:
  1. 1.
    $p(a,A)\geq 0$ for all $a\in X$
  2. 2.
    $p(a,A)=0$ if $a\notin A$
  3. 3.
    $\sum_{a\in A}p(a,A)\leq 1$
Below I abuse the notation to define what is called the default alternative, denoted by $a^{*}$. The default can be interpreted as abstention or failing to make a choice. The idea of default is pretty common in the literature on attention, especially the literature originating from Manzini and Mariotti (2014). The assumption is that DMs prefer to get something over abstention [^4] Equivalently, the default is only chosen if none of the alternatives is paid attention to.. For each $A\in\mathcal{X}$, define $p(a^{*},A)=1-\sum_{a\in A}p(a,A)$ to be probability of abstention from the menu $A$. I will also denote for any $A\in\mathcal{X}$, $A^{*}=A\cup\{a^{*}\}$.
The formal definition of a random attention model uses the notion of an attention function.
Definition 2
An attention function is a mapping $\mu:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$ such that:
  1. 1.
    $\sum_{D\subseteq A}\mu(D,A)=1$
  2. 2.
    $\mu(D,A)>0$ if and only if $D\subseteq A$
Here $\mu(D,A)$ can be interpreted as the probability that $D$ subset of $A$ is the consideration set when the menu $A$. It is natural that if $D$ is not a subset of $A$, then the function should take a value of zero. The second point also stresses the fact that each subset of $A$ is supposed to be a consideration set with some positive probability (full support assumption). This definition is looser than Cattaneo et al. (2020) since these authors, in contrast to the model presented below, impose the condition of monotonicity on the attention function $\mu$. A random attention model (RAM) is then defined as follows.
Definition 3
A stochastic choice function $p$ is a Random Attention Model (RAM) if there exists an attention function $\mu:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$ and a strict total order $\succ$ [^5] A strict total order is an asymmetric, transitive and complete binary relation defined over $X$. over $X$ such that for all $A\in\mathcal{X}$ and $x\in A$:
$p(x,A)=\sum_{D\in\{D\in\mathcal{X}:x\in D,x\succ y,\forall y\in D\setminus\{x\}\}}\mu(D,A)$
In other words, a random attention model is a stochastic choice function where for each menu, possible subsets of a menu (consideration sets) are considered/paid attention to with some positive probability. The probability that an alternative is chosen from a menu is the sum of the attention probabilities of consideration sets where that alternative is the most preferred with respect to the underlying preferences.
Since the empty set is also a possible consideration set for each menu, the probability of choosing the default is just the probability of the empty set being the consideration set. Note that RAM without the full support assumption has no empirical content since we can just attribute the probability of each alternative in a menu to the singleton menu containing that alternative and the probability of abstention to the attention probability of the empty set.
Now I will introduce the Random Attention Model with Unobserved References. Here I call the references to be exactly those alternatives that are attention privileged.
I start by first introducing a reference-dependent attention function, it being defined with respect to a pre-specified set $E\subseteq X$ of reference alternatives.
Definition 4
A reference-dependent attention function with the set of references being $E$ is a mapping $\mu_{E}:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$ such that:
  1. 1.
    $\sum_{D\subseteq A}\mu_{E}(D,A)=1$.
  2. 2.
    $\mu_{E}(D,A)>0$ if and only if $A\cap E\subseteq D\subseteq A$ [^6] This is the modified full support assumption, every subset which maintains the attention privilege of the references is paid attention to with positive probability..
The idea here is similar to attention privilege as defined by Kovach and Suleymanov (2023), i.e., if a reference alternative is in the menu, it is always considered. In fact, that is exactly the definition of a reference alternative in this setting. In my model, multiple reference alternatives or no reference alternatives can be present in any menu, both of which are not possible in the model presented by these authors since their primitives are menu-reference pairs. Note that the default is only chosen in case none of the alternatives are paid attention to, i.e. if $\mu_{E}(\phi,A)>0$. This can only happen if $A\cap E=\phi$. This allows us to interpret the default as a reference as well.
The fundamental analytical task in my model is to conduct a revealed preference exercise, i.e. identify the set of reference alternatives and the underlying preferences. In most canonical papers, the primitives assume observability of the references, which precludes the need to identify the set of references.
Definition 5
A stochastic choice function $p$ is a Random Attention Model with unobserved references (RAM-UR) if there exists a set $E\subseteq X$, a strict total order $\succ$ over $X$ and a reference-dependent attention function $\mu_{E}:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$ such that for all $A\in\mathcal{X}$ and $x\in A$:
$p(x,A)=\sum_{D\in\{D\in\mathcal{X}:x\succ y,\forall y\in D\setminus\{x\}\}}\mu_{E}(D,A)$
The above definition differs from Definition 3 in that it allows for only a reference-dependent attention function. This is the RAM with unobserved references in its most general form.
2.2 Characterization
I characterize the RAM-UR by utilizing three testable axioms. When we are dealing with stochastic choice functions axioms usually restrict how the choices behave across different menus.
Existence of a Dominant Alternative (EDA): For each $A\in\mathcal{X}$: $p(x,A)=0$ for some $x\in A^{*}$ implies $p(y,\{x,y\})=1$ for some $y\in A\setminus\{x\}$.
The idea captured by EDA is that of full support. The random process that determines the attention paid to various alternatives allows all alternatives to be considered with positive probability, so the only way an alternative can be never chosen from a menu is if whenever it is considered, it is dominated by another alternative which is considered. An alternative that is attention-privileged and preferred to it will certainly dominate it. It is also a necessary condition because of two properties that the full support assumption implies. First, it is not possible that an alternative is never chosen because it is never considered since all possible supersets of $A\cap E$ have a positive probability. Secondly, if there is no dominant alternative in the menu, there will be a subset of the menu with positive attention probability where the alternative will be the best.
Certainty WARP (C-WARP): For all $x\in X^{*}$ and $y\in X$ with $\{x,y\}\subseteq A^{*}\subseteq X^{*}$: $p(y,A)=1$ implies $p(x,A^{\prime})=0$ for all $A^{\prime}\subseteq X$ with $\{x,y\}\subseteq A^{\prime*}\subseteq X^{*}$.
Certainty WARP captures the notion that if an alternative is revealed to be dominated by an attention-privileged alternative, then it will never be chosen with a positive probability whenever that alternative is present. I call it C-WARP because it captures such revealed preference only when an alternative is chosen with certainty from a menu.
Expansion (EXP): For all $A,B\subseteq X$: $p(x,A)=p(y,B)=1$ for some $x\in A$ and $y\in B$ implies either $p(x,A\cup B)=1$ or $p(y,A\cup B)=1$.
Expansion captures the idea that as long as we keep observing certain choices in menus, there exists an attention-privileged alternative that dominates each alternative in the add of these menus. If we look at two attention-privileged alternatives, they will both be always considered, so the preference will determine which will be chosen, but it will be chosen with certainty in the union. It is in some ways similar to the expansion condition of Manzini and Mariotti (2007) in that it deals with the union of two menus. However, it works with any two menus with certain choices. It is silent on menus where certain choice is not observed.
My first result stated below shows that the above three axioms do allow a stochastic choice function to be represented by a RAM-UR model. However, the representation might not be unique.
Theorem 1
A stochastic choice function is representable by a random attention model with unobservable references if and only if it satisfies EDA, C-WARP and EXP.
It is worth mentioning that one can reach multiple representations of a stochastic choice function by slightly changing the constructed attention function. However, in each representation, the set of reference alternatives will be uniquely identified.
Even with limited data about choices from menus, in most cases we can identify the set of references uniquely. For instance, if there is no data available on singleton sets but we have the choice probabilities from all two and three-element menus, identification will be mostly possible. A rare exception is the case where the second least preferred alternative is not a reference but the least preferred alternative is. In such a case, we will not be able to ascertain which of the two will be the least preferred (and hence, a reference) and which is the second least preferred. However, for other alternatives, we will be able to identify uniquely.
The most important identification result is that we are only able to identify the preferences coarsely, i.e. we can identify the relative ranking of an alternative (whether a reference or not) with respect to all reference alternatives and the transitive closure of the partial order constructed through this relative ranking. However, for any two (non-reference) alternatives that have the same relative rankings with respect to all the reference alternatives, we are not able to identify the preference and thus can represent the choice function by taking any extension (Szpilrajn (1930)) of the revealed strict partial order [^7] A strict partial order is transitive, irreflexive and asymmetric binary relation defined over $X$.. The coarse identification is the result of the restrictions on both the observability of the references and on the amount of data available.
Kovach and Suleymanov (2023) were able to identify the complete underlying preferences from the general model with observed references. The RAM-UR works with stochastic choice functions and the references are unobservable. Given the significant restrictions to the primitives in this model it is natural that the preference identification result would be weaker. However, even in the restricted model, quite a significant part of the preferences are identified. Even among non-reference alternatives, such that they don’t have the same relative ranking with respect to the references, unique identification of the preferences among them is achieved.
A natural way to provide a full identification of preferences is to impose some restrictions on the structure of the attention function. In doing so, I characterize the RAM-UR-IRA by restricting the attention function to be of the Independent Random Attention (IRA) form as in Manzini and Mariotti (2014).
3 Random Attention Model with Unobserved References: Independent Random Attention
In this section I assume that the attention function in the RAM-UR model is of an independent random type in the sense that there exists an attention parameter $\gamma:X\rightarrow(0,1]$ such that the attention function takes the form: $\mu_{IRA}(D,A)=\prod_{x\in D\subseteq A}\gamma(x)\prod_{y\in A\setminus D}(1-(\gamma(y))$ for all $A\in\mathcal{X}$. Here attention privilege manifests in the form of alternatives with the attention parameter to be one, which is the key difference between this form and the attention function of the IRA form introduced by Manzini and Mariotti (2014) (hereafter referred to as the IRA model). The IRA restricts the attention parameter to be strictly between zero and one, we do not have such a restriction. This makes the IRA model to be nested in the RAM-UR-IRA.
Here the set of attention-privileged alternatives $E$ is just the set containing all alternatives $x\in X$ such that $\gamma(x)=1$. Since the $\gamma$ function takes strictly positive values for all alternatives, the RAM-UR-IRA is a special case of the RAM-UR as the full support assumption will hold.
Definition 6
A stochastic choice function $p$ is a Random Attention Model with unobservable references and independent random attention (RAM-UR-IRA) if there exists a strict total order $\succ$ over $X$ and an attention parameter $\gamma:X\rightarrow(0,1]$ such that for all $A\in\mathcal{X}$ and $x\in A$:
$p(x,A)=\gamma(x)\prod_{y\in A;y\succ x}(1-\gamma(y))$
Imposing the above additional structure significantly changes how the probabilities can be defined in the RAM-UR-IRA model. Even though it is a special case of RAM-UR and nests [^8] By nesting the IRA (rational) I mean that a special case of RAM-UR-IRA will also be an IRA (rational) model. both the IRA as well as classical rationality, it is mathematically a very small perturbation of the IRA model: It allows $\gamma$ to take the value one for some alternatives. I provide a characterization of this model by using axioms similar to the characterizations of both the IRA and the RAM-UR and discuss the identification results.
3.1 Characterization
I characterize the RAM-UR-IRA using the following five axioms:
R-Asymmetry (R-ASYM): For all $A\in\mathcal{X}$ and $a,b\in A$ with $p(a,A)>0$, $p(b,A)>0$: $\frac{p(a,A\setminus\{b\})}{p(a,A)}\neq 1$ implies $\frac{p(b,A\setminus\{a\})}{p(b,A)}=1$.
This axiom is similar to the i-asymmetry axiom that characterizes the IRA (cf. Manzini and Mariotti (2014)). However, it constrains the denominator to be non-zero. Such a constraint was not needed in the IRA characterization, since all probabilities there are assumed to be positive. This axiom says that if removing an alternative $b$ changes the choice probability of another, say $a$ (given the positive denominator constraints), then removing $a$ would not change the choice probability of $b$.
R-Independence (R-IND): For all $A,B\in\mathcal{X}$ and $a,b\in A\cap B$ with $p(a,A)>0,p(a,B)>0$: $\frac{p(a,A\setminus\{b\})}{p(a,A)}=\frac{p(a,B\setminus\{b\})}{p(a,B)}$.
For all $A,B\in\mathcal{X}$ and $b\in A\cap B$ with $p(a^{*},A)>0,p(a^{*},B)>0$: $\frac{p(a^{*},A\setminus\{b\})}{p(a^{*},A)}=\frac{p(a^{*},B\setminus\{b\})}{p(a^{*},B)}$.
This axiom is similar to the i-independence axiom in the characterization of the IRA (cf. Manzini and Mariotti (2014)), with the same constraints regarding non-zero probability when needed. It says that the ratio effect of removing an alternative on another alternative’s choice probability (subject to the positive denominator constraints), is menu-independent, i.e. it is the same whatever be the menu from which it is removed.
Non-Triviality (NT): $p(x,\{x\})>0$ for all $x\in X$.
This axiom implies that each alternative has a value on its own in the sense that if this axiom is not satisfied and we have an alternative that is not chosen in a singleton menu with positive probability, then other axioms will ensure that this alternative will not be chosen with a positive probability in any menu. Note that EDA introduced in Section 2.2 implies NT.
Existence of a Dominant Alternative* (EDA*): For all $A\in\mathcal{X}$: $p(a^{*},A)=0$, implies $p(a^{*},\{x\})=0$ for some $x\in A$.
This axiom is similar in spirit to the EDA axiom used in the characterization of RAM-UR. However, we only need EDA*, which is restricted to checking it for the default alternative, as opposed to checking for all alternatives that are chosen with zero probability in a menu. Note again that EDA implies EDA*, which means that NT and EDA* together are weaker than EDA.
Regularity (REG): For all $A,B\in\mathcal{X}$: $A\subset B$ implies $p(a,A)\geq p(a,B)$ for all $a\in A^{*}$.
This is the standard regularity axiom which states that an alternative’s choice probability weakly reduces in a larger menu. This property is in general not satisfied by the RAM-UR. If we consider the IRA model, the i-asymmetry and the i-independence axioms imply the regularity property. However, that is not the case with the other four axioms utilized to characterize the RAM-UR-IRA. The property is, thus, used as an axiom for the characterization. Note that this axiom implies that if we do not impose non-triviality, then any trivial alternative [^9] An alternative $x\in X$ here is called trivial if $p(x,\{x\})=0$. will not be chosen with a positive probability in any menu.
Theorem 2
A stochastic choice function is representable by a RAM-UR-IRA model if and only if it satisfies R-ASYM, R-IND, NT, EDA* and REG. Moreover, the representation is unique.
Representation uniqueness means that the set of references, the underlying preferences, and the underlying attention parameters are uniquely identified. This uniqueness result is due to the additional structure imposed on the attention function, and the availability of data in menus with no reference points.
The proof of Theorem 2 essentially separates the menus among those with no references and ones with at least one reference. Using the axioms above and the Manzini and Mariotti (2014) result on the set of menus with no reference, it identifies the underlying attention parameter and partial preferences among non reference alternatives (precisely the part of the preferences that were unidentified in the RAM-UR). Since the RAM-UR-IRA is a special case of the RAM-UR, the rest of the identification is easily done similar to the general model. The proof shows that the fitting of the identified preferences through the non-reference menus and the preferences identified through the menus with references does indeed give us a strict total order. Finally, using the first two axioms, we obtain the required functional form.
An advantage of a model like the RAM-UR is that it is in the realm of the stochastic choice function, which is why a nestedness analysis with some classical models can be undertaken. Both the RAM-UR and the RAM-UR-IRA nest rational choice as a special case, i.e. when the set of references is the entire set of alternatives. They both also nest the IRA, i.e. when the structure of the attention function is the IRA structure and the set of references is empty. Finally, the RAM-UR-IRA is also representable by a Random Utility Model (RUM) in a similar spirit to how the IRA can be. This is not true for the RAM-UR since it can violate regularity which is necessary for a RUM representation. A brief discussion on the RUM representation of the RAM-UR-IRA is provided in Appendix 2.
4 Literature Overview
The idea that limited attention affects behavior and choice is very natural and well-documented in the literature. Simon’s (1956) model of satisficing choice is one of the earliest models that can be motivated by limited attention, here the heuristic is that the DM stops at the first alternative which is “good enough”. A DM might utilize such a heuristic because of lack of time or attention. Lleras et al. (2017) model choice overload which leads to worse outcomes in larger menus, the reason being that good options might fail to grab attention in larger sets due to an overload of available options. Masatlioglu et al. (2012) introduce the concept of the attention filter which incorporates the idea that removing an unconsidered alternative does not change the consideration set in a deterministic choice setting where the DM is attention constrained. Dean et al. (2017) is another paper that incorporate limited attention and also the effect of reference alternatives.
The advent of stochastic choice as a fruitful direction of research to allow for the variability observed in choices called for modeling limited attention in such a framework. Manzini and Mariotti (2014) was the seminal paper that filled this void. Since then there has been plenty of literature combining both limited attention and stochastic choice. There are certainly lots of papers that are related closely with Manzini and Mariotti (2014), for instance, Brady and Rehbeck (2016)’s feasibility model is a generalization of the Manzini and Mariotti (2014), though their motivation is not limited attention, but limited feasibility and another one is Horan (2019) which looks at the IRA without a default. Cattaneo et al. (2020) introduces the general idea of a random attention model. Kovach and Suleymanov (2023) uses the idea of a RAM in the context of reference-dependent choice and modeled a family of attention functions, each indexed by the reference point, and looked at different special cases like the Constant Random Attention, Luce Random Attention and the Independent Random Attention. These authors discussed the nestedness properties among these special cases. They were not able to conduct the nestedness analysis with respect to the classical stochastic choice models because in their model the primitives were menu-reference pairs which put it outside the realm of stochastic choice models.
The idea of reference-dependent choice has been the subject of extensive study as well. Masatlioglu and Ok (2005) and Masatlioglu and Ok (2014) are seminal papers in this area. These authors were one of the first to model status quo bias in choice, which is a widely observed behavior in experiments on individual choice [^10] Masatlioglu and Ok (2005) provides a few sources for such experiments.. Sagi (2006) imposes a no-regret condition on the reference dependent preferences and Tapkı (2007) allows for incomplete reference dependent preferences. Note that all mentioned works model the effect of the status quo on preferences. However, it is possible that the status quo might affect attention as well. Dean et al. (2017) looks at status quo bias and limited attention together, allowing for both reference-dependent preferences and attention. Kovach and Suleymanov (2023), as discussed before, constructs a model where references affect only the attention. Bossert and Sprumont (2009) provides another model where reference points don’t affect preferences. They model a deterministic choice function where the choice is consistent with not choosing a worse alternative than the reference. They characterize where there exists a linear order such that the choice from each menu-reference pair is at least as good as the reference point, not necessarily the maximum [^11] The best alternative according to the preference ordering.. The RAM-UR does have this non-deteriorating property, since references are attention privileged, so in any menu an alternative worse than the reference is never chosen.
In the literature discussed above, a general theme is that the reference point is included as part of the primitive. It is natural to challenge the assumption of the observability of the reference point. Ok et al. (2015) and Tserenjigmid (2019) are two relevant papers here. However, in Tserenjigmid (2019) the references are not alternatives but a vector of attributes which depend upon the minimum of those attributes among the alternatives in any menu. It is thus, menu dependent as well. Kovach and Suleymanov (2023) is the first paper to study reference dependence in a stochastic setting, which makes my paper the first to study unobservable references in this framework.
We have assumed menu-independence with respect to the references. In Kovach and Suleymanov (2023), the input to the problem is menu-reference pairs so the question of menu independence was moot. However, it is possible to allow for menu-dependent unobservable references as a possible extension to the model in the current paper. This extension obviously has bite, i.e., we cannot explain any stochastic choice function through such a model. For example, any stochastic choice function that allows for alternative $x$ to be chosen with certainty in a menu where $y$ is present and allows alternative $y$ to be chosen with certainty in a menu where $x$ is present cannot be a reference-dependent RAM even with menu dependent references.
Appendix
Appendix 1
Proof of Theorem 1
Necessity:
EDA: Note that an alternative $x$ is never chosen from a menu if whenever it is paid attention to, a better alternative is paid attention to as well, which is only possible if a better alternative, say $y$ is attention privileged. In a binary menu including $x$ and $y$, $y$ will again be attention privileged so EDA will hold.
C-WARP: if $y$ is always chosen in a menu, then it is attention-privileged, and better than any other alternative because if it wasn’t, the full support assumption would mean that the other alternative would be chosen with positive probability. So, C-WARP will be satisfied.
EXP: If the antecedent of the axiom holds, then both $y$ and $x$ are attention-privileged, and every other alternative in $A^{*}$ and $B^{*}$ is less preferred to one of the two, so will not be chosen with positive probability. If $x$ is preferred to $y$ then it will always be chosen, and vice versa. So, EXP holds.
Sufficiency:
I take the set $\bar{E}=\{x\in X|p(x,\{x\})=1\}$ to be the revealed set of reference alternatives. If This set is empty EDA ensures that for any $A\in\mathcal{X}$, $p(a,A)>0$ for all $a\in A^{*}$. In this case, we can construct an attention function with $E=\bar{E}=\phi$ and an arbitrary preference. Note that with an empty set of references, the RAM-UR is just a RAM, so it suffices to show a RAM representation. For this, I take an arbitrary preference and construct an attention function as follows: For any alternative the probability that it is chosen is taken to be the sum of the value that the attention function takes on subsets where that alternative is the best. One way to do it is divide the probability of an arbitrary alternative $x$ in a menu $A$ to all sets in $\{D\subseteq A:x\succ y,\forall y\in D\setminus\{x\}\}$ equally ($\succ$ is the arbitrary preference). Note that since all alternatives (and default) are chosen with positive probability in all menus such a construction will give us a valid attention function. Given the constructed attention function, the observed choices are represented by a RAM with the attention function and the arbitrary preferences. Such a construction will be one possible RAM representation. We can have multiple representations if we change how we assign the probabilities. Thus, we have a representation when $\bar{E}$ is empty. For the rest of the proof I assume that $\bar{E}$ is non-empty.
In such a case, I define the binary relation $P$ over $X$ as follows:
  1. 1.
    For $x\in\bar{E}$ and $y\in X^{*}$ with $x\neq y$: $xPy$ iff $p(x,\{x,y\})=1$.
  2. 2.
    For $x\notin\bar{E}$ and $y\in\bar{E}$: $xPy$ iff $p(x,\{x,y\})>0$.
  3. 3.
    For $x,y\notin\bar{E}$: $xPy$ iff there exists a $z\in\bar{E}$ such that $xPz$ and $zPy$.
Lemma 1
$P$ is a strict partial order over $X$.
Proof. To show the asymmetry of $P$, we need to show that it is antisymmetric and irreflexive. I take it case-wise:
If $x\in\bar{E}$, then I haven’t defined $xPx$. If $x\neq y$, then $xPy$ implies $p(x,\{x,y\})=1$. So, $yPx$ is not possible since $p(y,\{x,y\})=0$, and $yPx$ cannot be the case. So for this case, there is no violation of irreflexivity and antisymmetricity.
If $x\notin\bar{E}$, then $xPx$ is only possible if there exists a $z\in\bar{E}$ such that $xPz$ and $zPx$ but that is not possible as above, so $P$ is irreflexive. Now, if $y\neq x$ such that $y\in\bar{E}$, then $xPy$ implies $p(x,\{x,y\})>0$, and thus, $yPx$ is not possible. Finally, if $y\neq x$ such that $y\notin\bar{E}$, then $xPy$ implies the existence of $z\in\bar{E}$ such that $p(x,\{x,z\})>0$ and $p(z,\{z,y\})=1$. If it were possible that there exists $w\in\bar{E}$ such that $p(w,\{w,x\})=1$ and $p(y,\{w,y\})>0$, then by EXP we have either $p(w,\{w,x,y,z\})=1$ or $p(z,\{w,x,y,z\})=1$ in contradiction to C-WARP. So, $P$ is asymmetric.
Now I show the transitivity of $P$. I will again argue casewise, in each of the cases having $xPy$ and $yPz$ for $x,y,z\in X$:
  • Case 1: $x\in\bar{E}$, $y\in\bar{E}$ and $z\in X^{*}$.
    By the definition of $P$ we have $p(x,\{x,y\})=1$ and $p(y,\{y,z\})=1$. EXP and C-WARP give us $p(x,\{x,y,z\})=1$, applying C-WARP again gives us $p(x,\{x,z\})=1$, so we have $xPz$.
  • Case 2: $x\in\bar{E}$, $y\notin\bar{E}$ and $z\in\bar{E}$.
    We have $p(x,\{x,y\})=1$ and $p(y,\{y,z\})>0$. Note that $p(z,\{z\})=1$ since $z\in\bar{E}$. So, by EXP, either $p(x,\{x,y,z\})=1$ or $p(z,\{x,y,z\})=1$, but the latter along with $p(y,\{y,z\})>0$ violates C-WARP, so we must have $p(x,\{x,y,z\})=1$ which by C-WARP gives us $p(x,\{x,z\})=1$ and thus, $xPz$.
  • Case 3: $x\in\bar{E}$, $y\notin\bar{E}$ and $z\notin\bar{E}$.
    We have $p(x,\{x,y\})=1$, as well as the existence of $w\in\bar{E}$ such that $yPw$ and $wPz$. Note that $w\neq x$ since $P$ is asymmetric. Looking at the menu $\{x,w,z\}$ we have $p(x,\{x,w,z\})=1$ because by Case 2 we know that $xPw$ and by EXP we know that either $p(x,\{x,w,z\})=1$ or $p(w,\{x,w,z\})=1$ holds. By C-WARP, $xPz$ follows.
  • Case 4: $x\notin\bar{E}$, $y\in\bar{E}$ and $z\in\bar{E}$.
    We have $p(x,\{x,y\})>0$ and $p(y,\{y,z\})=1$. If $p(x,\{x,z\})=0$, then we have $p(z,\{x,z\})=1$ by EDA. By EXP, either $p(y,\{x,y,z\})=1$ or $p(z,\{x,y,z\})=1$ must be the case, so $p(x,\{x,y,z\})=0$. By C-WARP, it must be that $p(z,\{x,y,z\})=1$, and applying C-WARP again we get $p(z,\{y,z\})=1$ in contradiction to $yPz$. So, it must be that $p(x,\{x,z\})>0$ and thus, $xPz$ follows.
  • Case 5: $x\notin\bar{E}$, $y\in\bar{E}$ and $z\notin\bar{E}$.
    We have $xPy$, $yPz$ and $y\in\bar{E}$, so by definition $xPz$ follows.
  • Case 6: $x\notin\bar{E}$, $y\notin\bar{E}$ and $z\in\bar{E}$.
    We have $xPy$ and thus, there exists $w\in\bar{E}$ such that $xPw$ and $wPy$. By Case 2 we have $wPz$, and by Case 4 we have $xPz$.
  • Case 7: $x\notin\bar{E}$, $y\notin\bar{E}$ and $z\notin\bar{E}$.
    We have $xPy$ and thus, there exists $w\in\bar{E}$ such that $xPw$ and $wPy$. So, by Case 3 we have $wPz$. By Case 5 $xPz$ follows.
In other words, the binary relation $P$ defined above is transitive. Note that P is complete when it comes to comparisons between an alternative in $\bar{E}$ and any other alternative. The only comparisons it could be silent on are between alternatives not in $\bar{E}$.
Given that $P$ is a strict partial order we can extend it to a strict total order [^12] A strict total order is an asymmetric, transitive and complete binary relation over $X$. (Szpilrajn (1930)). Note that such an extension need not be unique. Fix $\succ^{*}$ to be such an extension, that is, for all $x,y\in X$ we have that:
$xPy\implies x\succ^{*}y$
Next, I define the set $D_{x}$ to be the largest subset of $X$ such that $x$ is the $\succ^{*}$-maximal element.
Now I will specify the revealed attention function $\mu^{*}_{\bar{E}}:\mathcal{X}\times\mathcal{X}$ such that:
$p(x,A)=\sum_{D\in\{D:[x\cup(A\cap\bar{E})]\subseteq D\subseteq D_{x}\cap A\}}\mu^{*}_{\bar{E}}(D,A)$
We can have multiple attention functions that satisfy this requirement, I will show one way we can construct a revealed attention function.
Let $y$ be the $\succ^{*}$-maximal element in $A\cap\bar{E}$. Then any alternative $x\in A$ such that $y\succ^{*}x$ will have zero probability because of C-WARP. The set of possible sets $D$ that satisfy the requirement, $\{D:[x\cup(A\cap\bar{E})]\subseteq D\subseteq D_{x}\cap A\}$ is empty, so any attention function $\mu^{*}_{\bar{E}}$ will satisfy the above equality.
For an arbitrary element $x\in A$ which is $\succ^{*}$-better than $y$, EDA and C-WARP say that they will have positive probability in the menu. Similar to the construction by Kovach and Suleymanov (2023) we can ensure the existence of one such attention function by the following procedure: Let $n(x,A):=|\{D:[x\cup(A\cap\bar{E})]\subseteq D\subseteq D_{x}\cap A\}|$. Then we can specify $\mu^{*}_{\bar{E}}(D,A)=\frac{p(x,A)}{n(x,A)}$ for all $D\in\{D:[x\cup(A\cap\bar{E})]\subseteq D\subseteq D_{x}\cap A\}$. This gives us the required functional form. The set $\bar{E}$ is the revealed set of reference points which is unique. $\succ^{*}$ is the identified preferences unique up to the strict partial order $P$. $P$ uniquely gives us the preference of each alternative with respect to each reference points and the transitive closures that it implies.
Proof of Theorem 2
Necessity:
R-ASYM: If there exist $A\in\mathcal{X}$ and $x,y\in A$ with $p(x,A)>0$ and $p(y,A)>0$, then either none of $x$ and $y$ are references, in which case the axiom is true because of the functional form, as proved in Manzini and Mariotti (2014), or exactly one of them is a reference, and the other is preferred to it. In the latter case, the antecedent is true for the reference, since the probability of choice of the less preferred alternative (the reference) will increase if the other is removed, but removing the reference will not affect the choice probability of the better alternative.
R-IND: If there exists $A,B\in\mathcal{X}$, $a,b\in A\cap B$ with $p(a,A)>0$ and $p(a,B)>0$, the functional form is such that the R-IND will hold. Same is true if instead of $a$ we consider the default $a^{*}$. The proof is identical to the proof of the necessity of i-independence in Manzini and Mariotti (2014).
NT: This axiom is trivially satisfied given the functional form of the RAM-UR-IRA.
EDA*: If $p(a^{*},A)=0$, for some $A\in\mathcal{X}$ then there exists at least one alternative $x\in A$ such that $\gamma(x)=1$. Then $p(a^{*},\{x\})=0$ follows.
REG: Consider arbitrary $A\subseteq B\subseteq X$, then the functional form for any $x\in A$ gives us that $p(x,A)=\gamma(x)\prod_{y\in A;y\succ x}(1-\gamma(y))$. Similarly we get $p(x,B)=\gamma(x)\prod_{y\in B;y\succ x}(1-\gamma(y))$ for the set $B$. We can rewrite this as: $p(x,B)=\gamma(x)\prod_{y\in A;y\succ x}(1-\gamma(y))\prod_{y\in B\setminus A;y\succ x}(1-\gamma(y))$. Since all the attention parameters lie between zero and one, it is clear that $p(x,B)\leq p(x,A)$. So, REG is fulfilled.
Sufficiency:
I take the set $\bar{E}=\{x\in X:p(x,A)=1\hskip 1.42262pt\text{for some}\hskip 1.42262ptA\in\mathcal{X}\}$ to be the revealed set of reference alternatives.
Again, this set could be empty. If this set is empty, then EDA* and NT ensure that $p(a,A)>0$ for all $A\in\mathcal{X}$ and $a\in A^{*}$. Now the R-ASYM and the R-IND will be equivalent to the i-ASYM and i-IND by Manzini and Mariotti (2014), which gives us the required equivalence with the IRA model [^13] An IRA model is a RAM-UR-IRA model with $E=\phi$.. For the remainder of the proof, I assume non-emptyness of $\bar{E}$.
I will use the notation IRA to denote any reference to Manzini and Mariotti (2014) in this proof. We define two more properties RIIA and EXP, which the above axioms together imply:
Reference-Independence of Irrelevant Alternatives (RIIA): For all $A\in\mathcal{X}$ and $a\in A$: $p(a,A)=0$ implies $p(b,A\setminus\{a\})=p(b,A)$ for all $b\in A\setminus\{a\}$.
Claim 1: If a stochastic choice function $p:X\times\mathcal{X}\rightarrow[0,1]$ satisfies REG, then it satisfies RIIA.
Proof: By REG, for any alternative $c\in A^{*}\setminus\{a\}$ such that $p(c,A)>0$, we have $p(c,A)\leq p(c,A\setminus\{a\})$ and since summing over $p(c,A)$ for all $c\in A^{*}$ gives us one, they must be equal.
Expansion (EXP): For all $A,B\in\mathcal{X}$, $x_{1}\in A,x_{2}\in B$: $p(x_{1},A)=p(x_{2},B)=1$ implies either $p(x_{1},A\cup B)=1$ or $p(x_{2},A\cup B)=1$.
Claim 2: If a stochastic choice function $p:X\times\mathcal{X}\rightarrow[0,1]$ satisfies R-ASYM, R-IND and REG, then it satisfies EXP.
Proof: By REG, $p(y,A\cup B)=0$ for all $y\in A\cup B)^{*}\setminus\{x_{1},x_{2}\}$. Suppose $p(x_{1},A\cup B)>0$ and $p(x_{2},A\cup B)>0$. Then by R-IND, we have $\frac{p(x_{1},A\cup B\setminus\{x_{2}\})}{p(x_{1},A\cup B)}=\frac{p(x_{1},\{x_{1}\})}{p(x_{1},\{x_{1},x_{2}\})}$ and similarly for $x_{2}$. But $p(x_{1},\{x_{1}\})=p(x_{2},\{x_{2}\})=1$ by REG. So, $\frac{p(x_{1},A\cup B\setminus\{x_{2}\})}{p(x_{1},A\cup B)}>1$ and $\frac{p(x_{2},A\cup B\setminus\{x_{1}\})}{p(x_{2},A\cup B)}>1$ which contradicts R-ASYM.
I proceed with the proof of Theorem 2 and start with the following two lemmas.
Lemma 2
$A\cap\bar{E}=\phi$ for some $A\in\mathcal{X}$ implies $p(a,A)>0$ for all $a\in A$.
Proof. Suppose not. Then there exist a menu $A\in\mathcal{X}$ with $A\cap\bar{E}=\phi$ and $a\in A$ such that $p(a,A)=0$.
By NT and RIIA there exists $x\in A$ such that $p(x,A)>0$.
By RIIA, $\frac{p(x,A\setminus\{a\})}{p(x,A)}=1$.
Repeatedly applying RIIA we can remove irrelevant alternatives [^14] An alternative $a\in A\in\mathcal{X}$ is irrelevant with respect to $A$ if $p(a,A)=0$. from A to reach the menu $A^{{}^{\prime}}=\{x:x\in A,p(x,A)>0\}$. Note that $p(x,A)=p(x,A^{{}^{\prime}})$ for all $x\in A^{{}^{\prime}}$.
Also note that since $A^{\prime}\subseteq A$, $A^{\prime}\cap\bar{E}=\phi$. EDA* implies then $p(a^{*},A^{\prime})>0$. Due to the summation of choice probabilities, we have $p(a^{*},A^{\prime})=p(a^{*},A)$. Since $a\in A$ is such that $p(a,A)=0$ we have $p(a^{*},A)=p(a^{*},A\setminus a)$
Using R-IND and NT, we have $11=\frac{p(a^{*},A\setminus\{a\})}{p(a^{*},A)}=\frac{p(a^{*},\phi)}{p(a^{*},a)}>$ 1
which is a contradiction.
Denote by $\mathcal{X}^{{}^{\prime}}$ the set of menus which contain only the elements not in $\bar{E}$. I will first look only at the menus in the subsystem $\mathcal{X}^{{}^{\prime}}$. Lemma 2 indicates that for any menu $A\in\mathcal{X}^{{}^{\prime}}$, $p(a,A)>0$ for all $a\in A^{*}$.
Moreover, R-ASYM and R-IND reduce down to the i-asymmetry and the i-independence condition in the characterization of IRA, so in the subsystem $\mathcal{X}^{{}^{\prime}}$, the required functional form equivalence is achieved due to the characterization result by Manzini and Mariotti (2014). Also, the preferences are elicited using the i-regularity property which is also satisfied over this subsystem. Let the preferences elicited in such a way over alternatives in $X^{{}^{\prime}}$ be $\succ^{{}^{\prime}}$. Obviously, over this set the preference is a strict linear order.
Lemma 3
$A\cap\bar{E}\neq\phi$ for some $A\in\mathcal{X}$ implies $p(a^{*},A)=0$.
Proof. Suppose not. Then there exists $A\in\mathcal{X}$ such that $A\cap\bar{E}\neq\phi$ and $p(a^{*},A)>0$. For $x\in A\cap\bar{E}$, we have by REG that $p(a^{*},\{x\})>0$ or $p(x,\{x\})\neq 1$. Since $x\in\bar{E}$, $p(x,B)=1$ for some $B\in\mathcal{X}$. Applying REG leads to $p(x,\{x\})=1$ which is a contradiction.
I will now construct a binary relation over alternatives in $X$ and combine it with the preference $\succ^{{}^{\prime}}$ generated over $X^{{}^{\prime}}$ to form a preference over all alternatives. I will then show that this combination is a strict total order over all alternatives in $X$. For this, let us define the binary relation $\succ^{{}^{\prime\prime}}$ over $X$ as follows:
For all $a\in\bar{E}$ and $b\in X$: $p(a,\{a,b\})=1$ implies $a\succ^{{}^{\prime\prime}}b$, and $p(a,\{a,b\})\neq 1$ implies $b\succ^{{}^{\prime\prime}}a$.
I can put these two binary relations together to get another binary relation: $\succ=\succ^{{}^{\prime}}\cup\succ^{{}^{\prime\prime}}$.
Claim 3: The binary relation $\succ$ is a strict total order over $X$.
Proof. Obviously, $\succ$ is complete by definition. Among the alternatives in $X^{{}^{\prime}}$ it is clearly asymmetrical as well. We need to show asymmetry among alternatives where at least one of the alternatives is in $\bar{E}$.
Consider an arbitrary $x\in\bar{E}$ and arbitrary $y\in\bar{E}$ preferences are only defined through $\succ^{{}^{\prime\prime}}$ and is asymmetric by definition.
Suppose $a\in\bar{E}$ and $b\in\bar{E}$. By EXP we have either $p(a,\{a,b\})=1$ or $p(b,\{a,b\})=1$. Without loss of generality, if $p(a,\{a,b\})=1$, then $a\succ b$ but $b\nsucc a$. So, $\succ$ is asymmetric.
Now we show that $\succ$ is transitive as well. For this let us look at the following cases.
Case 1: $a\notin\bar{E}$, $b\notin\bar{E}$, $c\notin\bar{E}$, $a\succ b$ and $b\succ c$.
The binary relation $\succ$ here is defined entirely by $\succ^{{}^{\prime}}$ which is transitive as shown in the characterization of IRA, so transitivity of $\succ$ follows.
Case 2: $a\in\bar{E}$, $b\in\bar{E}$, $c\in X$, $a\succ b$ and $b\succ c$.
We have $p(a,\{a,b\})=1$ and $p(b,\{b,c\})=1$. By EXP, either $p(a,\{a,b,c\})=1$ or $p(b,\{a,b,c\})=1$ holds.
Suppose, $p(b,\{a,b,c\})=1$. Then by REG, $p(b,\{a,b\})=1$ follows. However, this contradicts the fact that $a\succ b$ and thus, $p(a,\{a,b,c\})=1$ should hold.
By REG again, we have $p(a,\{a,c\})=1$ and thus, $a\succ c$.
Case 3: $a\in\bar{E}$, $b\notin\bar{E}$ and $c\in\bar{E}$, $a\succ b$ and $b\succ c$.
We have $p(a,\{a,b\})=1$ and $p(b,\{b,c\})>0$ (otherwise, $p(a^{*},\{b,c\})=1$ in contradiction to Lemma 3).
Since $c\in\bar{E}$, we have by REG that $p(c,\{c\})=1$ holds. Using EXP, we have either $p(a,\{a,b,c\})=1$ or $p(c,\{a,b,c\})=1$.
If $p(c,\{a,b,c\})=1$, then by REG we have $p(c,\{b,c\})=1$ which is not true by assumption. So, $p(a,\{a,b,c\})=1$ should be the case.
By REG, we have $p(a,\{a,c\})=1$ and thus, $a\succ c$ follows.
Case 4: $a\in\bar{E}$, $b\notin\bar{E}$ and $c\notin\bar{E}$, $a\succ b$ and $b\succ c$.
We have $p(a,\{a,b\})=1$ and $\frac{p(b,\{b\})}{p(b,\{b,c\})}=1$ (which holds due to $i-asymmetry^{*}$ [^15] This condition is defined by Manzini and Mariotti (2014) which says that the implication in i-asymmetry goes both ways. They show that it is implied by their axioms. which holds over the menus $\{b\}$ and $\{b,c\}$ since they are part of the subsystem $\mathcal{X}^{{}^{\prime}}$ where all the IRA results apply).
Claim: $p(b,\{a,b,c\})=0$.
Proof. If not, then $p(b,\{a,b,c\})>0$. By REG, $p(b,\{a,b\})>0$ which is a contradiction.
By RIIA and the above claim $p(c,\{a,c\})=p(c,\{a,b,c\})$. Suppose $p(c,\{a,b,c\})>0$, then by R-IND, $1=\frac{p(c,\{a,c\})}{p(c,\{a,b,c\})}=\frac{p(c,\{c\})}{p(c,\{b,c\})}$ but $\frac{p(c,\{c\})}{p(c,\{b,c\})}\neq 1$ because of R-ASYM. Thus, it must be that $p(c,\{a,b,c\})=0$.
By RIIA, we have $p(c,\{a,c\})=p(c,\{a,b,c\})=0$. Applying Lemma 3, we get $p(a,\{a,c\})=1$. Thus, $a\succ c$ holds.
Case 5: $a\notin\bar{E}$, $b\notin\bar{E}$ and $c\in\bar{E}$, $a\succ b$ and $b\succ c$.
We have $\frac{p(b,\{b\})}{p(b,\{a,b\})}>1$ (by i-regularity of the IRA result on the subsystem $\mathcal{X}^{{}^{\prime}}$) and $p(b,\{b,c\})>0$ (by Lemma 3).
Claim: $p(a,\{a,b,c\})>0$.
Proof. Suppose not. Then $p(a,\{a,b,c\})=0$. By RIIA, $p(b,\{b,c\})=p(b,\{a,b,c\})>0$. By R-IND, $11=\frac{p(b,\{b,c\})}{p(b,\{a,b,c\})}=\frac{p(b,\{b\})}{p(b,\{a,b\})}>$ 1 which is not possible. So, $p(a,\{a,b,c\})>0$ follows.
By R-IND, we have $\frac{p(a,\{a,c\})}{p(a,\{a,b,c\})}=\frac{p(a,\{a\})}{p(a,\{a,b\})}$. We can use R-IND since by Lemma 1 we have $p(a,\{a,b\})>0$ as well. Note that by NT, we have $p(a,\{a\})>0$ and thus, we must have $p(a,\{a,c\})>0$. Hence, $a\succ c$ follows.
Case 6: $a\notin\bar{E}$, $b\in\bar{E}$ and $c\in\bar{E}$, $a\succ b$ and $b\succ c$.
We have $p(a,\{a,b\})>0$ (by Lemma 3) and $p(b,\{b,c\})=1$.
Claim: $p(a,\{a,b,c\})>0$.
Proof. Suppose not. then $p(a,\{a,b,c\})=0$. By RIIA we have $p(c,\{a,b,c\})=p(c,\{b,c\})=0$ and $p(b,\{a,b,c\})=p(b,\{b,c\})=1$. Applying RIIA again gives us $p(b,\{a,b\})=p(b,\{a,b,c\})=1$ which is not possible, so $p(a,\{a,b,c\})>0$.
By R-IND, $\frac{p(a,\{a,c\})}{p(a,\{a,b,c\})}=\frac{p(a,\{a\})}{p(a,\{a,b\})}$. This implies $p(a,\{a,c\})>0$ because $p(a,\{a\})>0$ and $p(a,\{a,b\})>0$. Thus, $a\succ c$ holds.
Case 7: $a\notin\bar{E}$, $b\in\bar{E}$ and $c\notin\bar{E}$, $a\succ b$ and $b\succ c$.
We have $p(a,\{a,b\})>0$ (by Lemma 3) and $p(b,\{b,c\})=1$.
Claim: $p(a,\{a,b,c\})>0$.
Proof. Suppose not. then $p(a,\{a,b,c\})=0$. By RIIA we have $p(c,\{a,b,c\})=p(c,\{b,c\})=0$ and $p(b,\{a,b,c\})=p(b,\{b,c\})=1$. Applying RIIA again gives us $p(b,\{a,b\})=p(b,\{a,b,c\})=1$ which is not possible, so $p(a,\{a,b,c\})>0$.
Note that $p(c,\{a,b,c\})=0$ by REG since $p(c,\{b,c\}=0$ by assumption.
Since $a\notin\bar{E}$, $p(a,\{a,b,c\})<1$. Since $p(c,\{a,b,c\})=0$ and $p(a^{*},\{a,b,c\})=0$ (by Lemma 3), we have $p(b,\{a,b,c\})>0$.
It follows from $\frac{p(b,\{b,c\})}{p(b,\{a,b,c\}}>1$ and R-ASYM that $\frac{p(a,\{a,c\})}{p(a,\{a,b,c\})}=1$ holds. By R-IND, we have $\frac{p(a,\{a\})}{p(a,\{a,c\})}=\frac{p(a,\{a,c\})}{p(a,\{a,b,c\})}=1$. In the subsystem containing the menu $\{a,c\}$, $i-asymmetry^{*}$ holds and thus, $\frac{p(c,\{c\})}{p(c,\{a,c\})}>1$ leads to $a\succ c$.
In view of the above cases we have that $\succ$ is transitive and thus, a strict total order over $X$.
We now define $\gamma:X\rightarrow(0,1]$ by $\gamma(a)=p(a,\{a\})$ for each $a\in X$. Note that $\gamma(a)=1$ for all $a\in\bar{E}$ and $\gamma(b)\in(0,1)$ for all $b\notin\bar{E}$. $\bar{E}$ will be identified here as the set $E$ of the RAM-UR-IRA.
Claim 4: For all $A\in\mathcal{X}$ and $a,b\in A$: $p(a,A)>0$, $p(b,\{b\})\neq 1$ and $p(a,A\setminus\{b\})>p(a,A)$ imply $\frac{p(a,A\setminus\{b\})}{p(a,A)}=\frac{1}{1-p(b,\{b\})}$ for all $a\in A^{*}$.
Proof. This claim is similar to the one in the proof of the first theorem in Manzini and Mariotti (2014). Proving this claim will allow us to recover the functional form of the IRA structure.
For $a^{*}$, the required equality translates to $\frac{p(a^{*},A\setminus\{b\})}{p(a^{*},A)}=\frac{1}{1-p(b,\{b\})}$. Note that $\frac{1}{1-p(b,\{b\})}=\frac{p(a^{*},\phi)}{p(a^{*},\{b\})}$. which is true by the implementation of R-IND.
For $a,b\in A$ such that $\frac{p(a,A\setminus\{b\})}{p(a,A)}>1$. Here, RIIA implies that $p(b,A)>0$.
We can apply R-ASYM, to get $p(b,A\setminus\{a\})=p(b,A)$
By REG we get $p(a,\{a,b\})>0$.
By R-IND we have,
$\frac{p(a,A\setminus\{b\})}{p(a,A)}=\frac{p(a,\{a\})}{p(a,\{a,b\})}=\frac{1-p(a^{*},\{a\})}{1-p(b,\{a,b\})-p(a^{*},\{a,b\})}$
. Applying R-ASYM,
$\frac{p(a,A\setminus\{b\})}{p(a,A)}=\frac{1-p(a^{*},\{a\})}{1-p(b,\{b\})-p(a^{*},\{a,b\})}$
If $p(a,\{a\})=1$, then we get the required equation by REG. If not we have $a\notin\bar{E}$ so by EDA* we have $p(a^{*},\{a,b\})>0$. We can then use the following equation from the first part of this proof:
$p(a^{*},\{a\})=\frac{p(a^{*},\{a,b\})}{1-p(b,\{b\})}$
Putting this above we will get the required equality:
$\frac{p(a,A\setminus\{b\})}{p(a,A)}=\frac{1}{1-p(b,\{b\})}$
Now I will look at different set of menus and alternatives to show that we get the required functional form.
For $a\in\bar{E}$, define $A=\{a\}\cup\{x:x\in X,a\succ X\}$. By definition of $\succ$, $p(a,\{a,b\})=1$ for all $b\in A$. So, by REG, $p(a,A)=1=\delta(a)$. As required.
Let $A\in\mathcal{X}$ be a menu such that there exists atleast one $x\in A\cap\bar{E}$. Let $a\in A\cap\bar{E}$ such that $a\succ b$ for any $b\in A\cap\bar{E}\setminus\{a\}$, essentially, $a$ is the most preferred alternative in $A\cap\bar{E}$.
For any alternative $b\in A$ such that $a\succ b$, $p(b,A)=0$ by REG. Also by REG, $p(a^{*},A)=0$ as required.
Claim: In the above case, $p(a,A)>0$.
Proof. Suppose not. Then $p(a,A)=0$.
Since we have assumed $p(a,A)=0$, $p(c,A)=p(c,A\setminus\{a\})$ follows for all $c\in A\setminus\{a\}$. We can iteratively remove all alternatives in $A$ that are worse than $a$ and $a$ itself to reach $A^{{}^{\prime}}=\{b\in A:b\succ a\}$.
Since, $a$ is the most preferred alternative in $\bar{E}$ that is in $A$, $A^{{}^{\prime}}\cap\bar{E}=\phi.$. So by EDA*, $p(a^{*},A^{{}^{\prime}})>0$.
Since $p(b,A)=p(b,A^{{}^{\prime}})$ for all $b\in A^{{}^{\prime}}$ and $p(c,A)=0$ for all $c\in A\setminus(A^{{}^{\prime}})$, $p(a^{*},A)>0$ follows in contradiction to REG.
Claim: If $c\succ a$ and $c\in A$, then $p(c,A)>0$.
Proof. Suppose, there exists $c\in A$ such that $p(c,A)=0$ and $c\succ a$. Since $p(a,A)>0$, we have by R-IND $1<\frac{p(a,\{a\})}{p(a,\{a,c\})}=\frac{p(a,A\setminus\{c\})}{p(a,A)}$ which is incompatible with RIIA. So, it must be that $p(c,A)>0$ for all $c\succ a$.
I can now iteratively apply Claim 4 similar to how it was done in Manzini and Mariotti (2014) to get the required equivalence with the functional form. We have already considered menus without any alternatives in $\bar{E}$ through the subsystem $\mathcal{X}^{{}^{\prime}}$. The rest of the cases have been considered above. $E$ is identified uniquely as the $\bar{E}$ and preferences as specified above are identified uniquely by $\succ$.
Appendix 2
In this section, I discuss the Random Utility Model representation of the RAM-UR-IRA and analyze how the representation changes as the set of reference alternatives $E$ changes.
Denote by $\mathcal{L}(X^{*})$ the set of all strict total orders over $X^{*}$ with typical element $R$. A probability distribution over $\mathcal{L}(X^{*})$ is denoted by $\nu$.
Definition 7
A stochastic choice function $p:X^{*}\times\mathcal{X}$ has a Random Utility Model (RUM) representation if there exists a distribution $\nu$ over the set of all strict total orders on $X^{*}$, $\mathcal{L}(X^{*})$ such that:
$p(x,A)=\sum_{R\in\mathcal{L}(X^{*}):xRy\forall y\in A^{*}\setminus\{x\}}\nu(R)$
for all $x\in A^{*}$.
The representation is very similar in construction to the one provided by Manzini and Mariotti (2014) for their IRA model which is a RUM too. The construction for the IRA goes as follows:
  • For any alternative $a\in X$, set $\nu(\{R:aRa^{*}\})=\gamma(a)$.
  • If $a\succ b$, then for an $R$ such that $bRa$, $aRa^{*}$ and $bRa^{*}$, set $\nu(R)=0$.
  • For any two alternatives $a$ and $b$, $\nu(\{R:aRa^{*}\hskip 2.84526pt\text{and}\hskip 2.84526ptbRa^{*}\})=\gamma(a)\gamma(b)$.
The above restrictions continue to hold for the RAM-UR-IRA but we will also need more restrictions for our model:
  1. 1.
    A reference is always considered if available. So, we have $\nu(\{R:a^{*}Ra\})=0$ for all $a\in E$.
  2. 2.
    If $a\in E$ and $a\succ b$ for $b\in X$, set $\nu(\{R:b\succ a\})=0$.
(1) here is implied by the first point in the construction by Manzini and Mariotti (2014). It is reiterated to emphasize that all strict orders where the default beats a reference alternative are essentially removed from the support of the RUM.
Note that as we add more alternatives to the set $E$, more and more strict total orders are removed from the support. As $E$ becomes the set $X$, only the strict linear order, which is the preference of the agent is in the support, and the agent is rational. On the other hand, if $E$ is the empty set, the first set of restrictions will be the only ones that apply, and we are back to the Manzini and Mariotti (2014) model. Looking at the RUM representation makes the nestedness properties of the RAM-UR-IRA clearer.
Note finally that the RAM-UR might not be a RUM because it fails to satisfy regularity which is a necessary condition for a stochastic choice function to be a RUM, (Falmagne (1978)). This is the case because of the lack of structure on the attention function due to which for an alternative that is undominated by any reference in a smaller and a bigger menu, we can construct attention probabilities of the subsets where it is the best alternative to be more in the bigger menu than in the smaller menu. This will lead to a RAM-UR model that violate regularity. The lack of structure of the general model allows us to construct attention functions that gives us the means to explain different behavioral observations. One such empirically observed behavioral pattern that it allows us to explain is that the status quo (references) is more likely to be chosen in larger sets (Choice Overload) similar to how Kovach and Suleymanov (2023) showed in their model using the Luce Random Attention structure.