### v4

Abstract
In this paper, I develop and characterize a random attention model with unobserved reference alternatives. The decision-maker pays attention to different subsets of the available set of alternatives randomly. The reference alternatives are exactly those alternatives that are always paid attention to, i.e. they are attention-privileged. These alternatives are unknown to the outside observer. The characterization allows for a complete identification of the reference alternatives and a coarse identification of the underlying preferences. I then restrict the model by considering the independent random attention function and provide a complete identification of the underlying preferences.
JEL Classification: D01, D91, Z13
Keywords: Bounded Rationality, Stochastic Choice, Limited Attention, Consideration Sets
1 Introduction
Decision-making is often constrained by limited attention, whether due to cognitive processing limits caused by the availability of too many alternatives or a shortage of time. In such situations, decision-makers (DM) often resort to heuristics that can help them decide relatively quickly with less mental exercise. If the DM has to make similar choices multiple times with the same underlying primitives, the heuristic that decides how to allocate attention can lead to different choices. Stochastic choice functions represent such choices, with the main research task then being to ascertain the likely underlying heuristic and the primitives that drive the decisions of the DM. Cattaneo et al. (2020) introduced the Random Attention Model (RAM) framework to deal with the analysis of stochastic choice functions with random choice due to limited attention.
A substantial body of literature has explored how an observable reference affects the choices of a DM. In particular, Kovach and Suleymanov (2023) uses the framework of RAM to look at how a reference alternative, or a status quo in their case, can affect choices if it directs attention. They were the first to deal with observable reference alternatives’ impact on attention in a stochastic choice setting. It is reasonable to assume that a reference alternative can direct attention when the underlying heuristic is applied due to limited attention. They provide a characterization of a reference-dependent RAM, as well as special cases where the attention function is further restricted.
In certain cases, it is logical to assume that the reference is observable. For instance, for a DM in the market for a new car, her current car can be a reference. However, in many situations, it is not realistic to assume the knowledge of the DM’s reference to anyone else but the DM themselves. Further, to analyze a choice problem with observable references, we have to deal with a family of stochastic choice functions, as done by Kovach and Suleymanov (2023). The inputs for the stochastic choice functions in their model were pairs of subsets [^3] Hereafter referred to as menus. and references- one family of stochastic choice function for each alternative as the reference. This is quite a significant data requirement. If we restrict the references to be unknown, we deal with just a single stochastic choice function over menus and thus, we do not have to deal with families of stochastic choice functions for each possible reference.
It is also not obvious that a reference alternative must be unique. In what follows, I allow for multiple reference alternatives; clearly, one particular such alternative could be the status quo. In this spirit, I define an alternative to be a reference alternative if and only if it is attention privileged, i.e. it is always paid attention to. Since I am looking at choices from menus that depend on unobservable reference alternatives, I assume that the references are menu-independent. It is certainly plausible that references can be menu-dependent, through which we can explain a larger set of stochastic choice behavior, but dealing with menu-dependent unobservable references is beyond the scope of this paper. I provide below four examples that match the mentioned assumptions.
  • Example 1: University Choice: High School students making a choice of university to continue their higher education in the UK filter out the large number of choices into smaller consideration sets. The universities are divided into old and new universities. This exogenous distinction impacts what universities are considered by students. For instance, Dawes and Brown (2005) found that minority ethnic groups had a greater proportion of old universities in their consideration set, possibly due to reputation concerns of these ethnic groups. The college choice of high school students in the US is also a pretty involved process. Rosen et al. (2008) showed that what major is offered as well as location were crucial factors in shortlisting colleges. Parental and guidance counselor advice also has a significant impact on the decision-making process. It is conceivable that colleges that are close by act as references. Dawes and Brown (2005) observed a similar phenomenon in the UK data as well. If the guidance of a counselor or a parent impacts which colleges act as references, it is not possible to observe these references.
  • Example 2: Apparel Shopping: While shopping for apparel, brand name is often crucial in drawing attention to particular items. The Air Jordan always catches the eye due to its association with Michael Jordan. However, though brand awareness might be enough to pull attention, it is not enough to provide a positive brand association as found by Dew and Kwon (2009). Thus, attention and preference are independent attributes.
  • Example 3: Political Choice: A stochastic choice framework can also be used to model a population of voters with similar preferences but with access to different options. A nationally strong party (a reference) will always be considered; however, a weaker party in one district might be a stronger party in another and be considered there, leading to different choices even when voters exhibit homogenous preferences. Wilson (2008) showed this for the Mexico national elections in 2000. The biggest party in Mexico, the Partido Revolucionario Institucional (PRI) has a strong presence in all the states, which makes it the reference, while the two opposition parties have varied presence in different states.
  • Example $4$I would like to thank Marco Mariotti for this interpretation of the model.: Suppose past choices of the DM influence their future choices, i.e. they act as a reference. The outside observer only observes the current feasible sets and the choices from those but not the choices in the past. In this case, (multiple) past choices act as references and are thus, unobservable to the researcher. If the reference depends on recommendations from advisors (college choice) or past choices in elections (political choice), it is possible that the researcher is not able to observe the reference.
In this paper, I look at a random attention model with unobservable menu-independent reference alternatives that are attention-privileged (RAM-UR). I provide a characterization of the RAM-UR by using three axioms and identifying the corresponding primitives (i.e., the set of reference alternatives and the underlying preferences). One can uniquely identify the set of references but the underlying preferences are only coarsely identified. The identification of the preferences gives the relative ranking between any alternative and any reference alternative and the transitive closure thereof. I then restrict the attention function to be of the independent random attention (IRA) form (Manzini and Mariotti (2014)) and show that under such a restriction, the preferences can be fully identified. The existence of menus with no references is crucial in making the preference identification in this case. This is possible because of the menu-independence of the references. Kovach and Suleymanov (2023) allowed for the references to be part of the primitive which precludes them from considering menus without references. I also elaborate on the fact that rational choice is a boundary case of both the RAM-UR and the RAM-UR-IRA, while the IRA model is itself a boundary case of the RAM-UR-IRA [^5] Boundary cases can be of two types: When the set of references is the set of all alternatives, or when the set of references is empty..
The rest of the paper is organized as follows. In Section 2, I introduce the general RAM-UR model, the axiomatic characterization of the model is presented in Section 3 along with a discussion on the identification results. The introduction and characterization of the RAM-UR-IRA are then provided in Section 4, where I also discuss the nestedness properties of the corresponding random attention models. Section 5 relates my work to the existing literature. All proofs are relegated to Appendix 1.
2 Random Attention Model with Unobserved References
Let $X$ be the finite set of all possible alternatives. Denote the set of all subsets of $X$ by $\mathcal{X}$, with a typical element being $A$.
Definition 1
A stochastic choice function is a mapping $p:X\times\mathcal{X}\setminus\phi\rightarrow[0,1]$ where for each $A\in\mathcal{X}\setminus\phi$ the following three conditions hold:
  1. 1.
    $p(a,A)\geq 0$ for all $a\in X$
  2. 2.
    $p(a,A)=0$ if $a\in X\setminus A$
  3. 3.
    $\sum_{a\in A}p(a,A)\leq 1$
Below I denote by $a^{*}$ the default alternative and interpret it as abstention or failing to make a choice or delaying choice. The idea of default is common in the literature on attention, especially the literature originating from Manzini and Mariotti (2014). The assumption is that DMs prefer to get something over abstention [^6] Equivalently, the default is only chosen if none of the alternatives is paid attention to.. For each $A\in\mathcal{X}\setminus\phi$, define $p(a^{*},A)=1-\sum_{a\in A}p(a,A)$ to be probability of abstention from the menu $A$. I will also denote for any $A\in\mathcal{X}\setminus\phi$, $A^{*}=A\cup\{a^{*}\}$.
The formal definition of a random attention model uses the notion of an attention function.
Definition 2
An attention function is a mapping $\mu:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$ such that:
  1. 1.
    $\sum_{D\subseteq A}\mu(D,A)=1$
  2. 2.
    $\mu(D,A)>0$ if and only if $D\subseteq A$
Here $\mu(D,A)$ can be interpreted as the probability that $D$ is the consideration set when the menu offered is $A$. It is natural that if $D$ is not a subset of $A$, then the function should take a value of zero. The second point also stresses the fact that each subset of $A$ can be a consideration set with some positive probability (full support assumption). In contrast to the above definition, Cattaneo et al. (2020) impose a monotonicity condition on the attention function $\mu$ and do not impose the full support assumption. Kovach and Suleymanov (2023) imposed the full support assumption and it will be crucial in our model as well to identify the set of references. Without the full support assumption, we cannot differentiate between the reference and some of the non-reference alternatives. I also relax the monotonicity condition in the general model. As demonstrated later, this relaxation does not result in much loss in identification. A random attention model (RAM) is then defined as follows.
Definition 3
A stochastic choice function $p$ is a Random Attention Model (RAM) if there exists an attention function $\mu:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$ and a strict total order $\succ$ [^7] A strict total order is an asymmetric, transitive and complete binary relation defined over $X$. over $X$ such that for all $A\in\mathcal{X}\setminus\phi$ and $x\in A$:
$p(x,A)=\sum_{D\in\{D\in\mathcal{X}:x\in D,x\succ y,\forall y\in D\setminus\{x\}\}}\mu(D,A)$
In other words, a random attention model is a stochastic choice function where for each menu, each possible subset of a menu (consideration sets) is considered/paid attention to with some positive probability. The probability that an alternative is chosen from a menu is the sum of the attention probabilities of consideration sets where that alternative is the most preferred with respect to the underlying preferences.
Since the empty set is also a possible consideration set for each menu, the probability of choosing the default is just the probability of the empty set being the consideration set. Note that the RAM without the full support assumption has no empirical content since we can just attribute the probability of each alternative in a menu to the singleton menu containing that alternative and the probability of abstention to the attention probability of the empty set. The empirical content of the RAM in Cattaneo et al. (2020) comes from the monotonicity restriction they impose on the attention function.
Below I introduce the Random Attention Model with Unobserved References with references being exactly those alternatives that are attention privileged.
I start by first introducing the notion of a reference-dependent attention function.
Definition 4
Let $E\subseteq X$ be a set of reference alternatives. A reference-dependent attention function with respect to $E$ is a mapping $\mu_{E}:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$ such that:
  1. 1.
    $\sum_{D\subseteq A}\mu_{E}(D,A)=1$
  2. 2.
    $\mu_{E}(D,A)>0$ if and only if $A\cap E\subseteq D\subseteq A$ [^8] This is the modified full support assumption, every subset which maintains the attention privilege of the references is paid attention to with positive probability.
The idea here is similar to attention privilege as defined by Kovach and Suleymanov (2023), i.e., if a reference alternative is in the menu, it is always considered. In fact, that is exactly how I define a reference alternative in this setting. A reference alternative can have other properties, which maybe correlated with the preferences, dependent upon some status quo, or some attribute (for instance, the color of a car). However, the presented model is silent on any such dependence. An alternative is a reference if and only if it is attention privileged, irrespective of how it attained such a privilege. In my model, multiple reference alternatives or no reference alternatives can be present in any menu, both of which are not possible in the model presented by Kovach and Suleymanov (2023) since the primitives there are menu-reference pairs. Note that the default is only chosen in case none of the alternatives are paid attention to, i.e. if $\mu_{E}(\phi,A)>0$. This can only happen if $A\cap E=\phi$ and allows us to interpret the default as a reference.
The fundamental analytical task with respect to the presented model is to identify the set of reference alternatives and the underlying preferences. In most canonical papers, it is assumed that the references are observable, which precludes the need to identify the set of references.
Definition 5
A stochastic choice function $p$ is a random attention model with unobserved references (RAM-UR) if there exists a set $E\subseteq X$, a strict total order $\succ$ over $X$ and a reference-dependent attention function $\mu_{E}:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$ such that for all $A\in\mathcal{X}\setminus\phi$ and $x\in A$:
$p(x,A)=\sum_{D\in\{D\in\mathcal{X}:x\succ y,\forall y\in D\setminus\{x\}\}}\mu_{E}(D,A)$
The above definition differs from Definition 3 in that it allows for only a reference-dependent attention function. This is the RAM with unobserved references in its most general form.
3 Characterization
In the present section I characterize the RAM-UR by using three testable axioms. These axioms restrict how the stochastic choice function operates across different menus.
Existence of a Dominant Alternative (EDA): For each $A\in\mathcal{X}\setminus\phi$: $p(x,A)=0$ for some $x\in A^{*}$ implies $p(y,\{x,y\})=1$ for some $y\in A\setminus\{x\}$.
A DM with limited attention fails to choose an alternative in two cases. First, if that alternative is never considered. Second, if whenever it is considered, it is dominated by another alternative. EDA says that, the only way an alternative is never chosen is if another alternative always dominates it and is always considered in the binary choice between the two. This means that it is not possible for an alternative to never be chosen just because it is never considered, making each alternative to be non-trivial in the problem. The axiom also says that the set of possible consideration sets is rich enough that the only way an alternative is never chosen is if another dominating alternative is always considered, i.e. it is attention privileged. This way, EDA captures the essence of the modified full support assumption.
Certainty WARP (C-WARP): For all $x\in X^{*}$ and $y\in X$ with $\{x,y\}\subseteq A^{*}\subseteq X^{*}$: $p(y,A)=1$ implies $p(x,A^{\prime})=0$ for all $A^{\prime}\subseteq X$ with $\{x,y\}\subseteq A^{\prime*}\subseteq X^{*}$.
Certainty WARP captures the idea that if an alternative is revealed to be dominated by an attention-privileged alternative, then it will never be chosen with a positive probability whenever that alternative is present. I call it C-WARP because it captures such revealed preference only when an alternative is chosen with certainty from a menu.
Expansion (EXP): For all $A,B\subseteq X$: $p(x,A)=p(y,B)=1$ for some $x\in A$ and $y\in B$ implies either $p(x,A\cup B)=1$ or $p(y,A\cup B)=1$.
Expansion captures the idea that as long as we keep observing certain choices in menus, there exists an attention-privileged alternative that dominates each alternative in the union of these menus. If we look at two attention-privileged alternatives, they will both be always considered, so the preference will determine which will be chosen, but it will be chosen with certainty in the union. It is in some ways similar to the expansion condition of Manzini and Mariotti (2007) in that it deals with the union of two menus. However, it works with any two menus with certain choices. It is silent on menus where certain choice is not observed.
The three axioms are independent and restrict the stochastic functions only where deterministic choice or non-choice is observed. This is because the model is very flexible on the structure of the attention function, once the full support assumption is satisfied. To further restrict the stochastic function some regularity condition similar to Cattaneo et al. (2020) must be imposed. However, such an imposition is not necessary to get the first set of results described below.
My first result shows that the above three axioms do allow a stochastic choice function to be represented by a RAM-UR model. However, the representation might not be unique.
Theorem 1
A stochastic choice function is representable by a random attention model with unobservable references if and only if it satisfies EDA, C-WARP and EXP.
It is worth mentioning that one can reach multiple representations of a stochastic choice function by slightly changing the constructed attention function. However, in each representation, the set of reference alternatives will be uniquely identified.
Even with limited data about choices from menus, in most cases we can identify the set of references uniquely. For instance, if there is no data available on singleton sets but we have the choice probabilities from all two and three-element menus, identification will be mostly possible. A rare exception is the case where the second least preferred alternative is not a reference but the least preferred alternative is. In such a case, we will not be able to ascertain which of the two will be the least preferred (and hence, a reference) and which is the second least preferred. However, we will be able to identify all other reference alternatives uniquely.
The most important identification result is that we are only able to identify the preferences coarsely, i.e. we can identify the relative ranking of an alternative (whether a reference or not) with respect to all reference alternatives and the transitive closure of the partial order constructed through this relative ranking. However, for any two (non-reference) alternatives that have the same relative rankings with respect to all the reference alternatives, we are not able to identify the preference and thus, one can represent the choice function by taking any extension (Szpilrajn (1930)) of the revealed strict partial order [^9] A strict partial order is transitive, irreflexive and asymmetric binary relation defined over $X$.. The coarse identification is the result of the restrictions on both the observability of the references and on the amount of data available.
Even though the axioms restrict choice only where deterministic choice or non-choice is observed, it is not necessary to observe certain choice to identify the set of references. It is also not necessary to observe all menus to make such an identification. The following example illustrates that:
Example 5
Suppose we have four alternatives $X=\{a,b,c,d\}$ and the following observed choice probabilities:
$p(a,\{a,b,c\})=0.4\hskip 14.22636ptp(b,\{a,b,c\})=0.3\hskip 14.22636ptp(c,\{a,b,c\})=0.3$
$p(a,\{a,c,d\})=0.6\hskip 14.22636ptp(c,\{a,c,d\})=0.4\hskip 14.22636ptp(d,\{a,c,d\})=0$
$p(a,\{a,b,d\})=0.4\hskip 14.22636ptp(b,\{a,b,d\})=0.3\hskip 14.22636ptp(d,\{a,b,d\})=0.2$
$p(b,\{b,c,d\})=0.5\hskip 14.22636ptp(c,\{b,c,d\})=0.5\hskip 14.22636ptp(d,\{b,c,d\})=0$
Here we observe data only from menus of size three, and we do not observe any deterministic choice.
Note that in the above example, we can identify alternative $c$ as the only reference, as well as the partial order without the need for the complete set of menus. In fact, in any setting with a large number of alternatives and relatively few references, we can make full identification with very few data points, and we will not need to observe a lot of deterministic choices in those menus.
Kovach and Suleymanov (2023) were able to identify the complete underlying preferences from the general model with observed references. The RAM-UR works with stochastic choice functions and the references are unobservable. Given the significant restrictions to the primitives in this model, it is natural that the preference identification result is weaker. However, even in the restricted model, quite a significant part of the preferences are identified. Even among non-reference alternatives, with different relative ranking with respect to the references, unique identification of the preferences among them is achieved.
A natural way to provide a full identification of preferences is to impose some restrictions on the structure of the attention function. First, I will show that the monotonicity condition imposed by Cattaneo et al. (2020) does not offer us any progress in the identification results.
Definition 6
A reference-dependent monotonic attention function with respect to a set $E\subseteq X$ of references is a mapping $\mu_{E}:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$ such that:
  1. 1.
    $\sum_{D\subseteq A}\mu_{E}(D,A)=1$
  2. 2.
    $\mu_{E}(D,A)>0$ if and only if $A\cap E\subseteq D\subseteq A$
  3. 3.
    For $S,T\in\mathcal{X}$, if $a\in S\setminus T$, then $\mu_{E}(T,S\setminus\{a\})\geq\mu_{E}(T,S)$
The example below illustrates that even if we restrict ourselves to monotonic attention functions, a unique identification of the preferences is not always possible.
Example 6
Suppose we have three alternatives $X=\{a,b,c\}$ and the following observed choice probabilities:
$p(b,\{a,b\})=1/2\hskip 14.22636ptp(c,\{a,c\})=1/2\hskip 14.22636ptp(a,\{a,b\})=p(a,\{a,c\})=1/2$
$p(a,\{a,b,c\})=p(c\{a,b,c\})=1/4\hskip 14.22636ptp(b,\{a,b,c\})=1/2$
$p(a,\{a\})=1\hskip 14.22636ptp(b,\{b\})=p(c,\{c\})=1/2$
$p(b,\{b,c\})=1/4\hskip 14.22636ptp(c,\{b,c\})=1/2$
We can rationalize these observed choice probabilities through two reference-dependent monotonic attention functions. In both cases $E=\{a\}$ and both $b$ and $c$ are revealed preferred to $a$.
  • $\mu^{1}_{E}$ is the function such that attention is divided among all eligible [^10] Supersets of {a} in any menu containing $a$, and all subsets in any other menu. sets in each menu with the preference being $b\succ c\succ a$.
  • Construct $\mu^{2}_{E}$ as follows:
    $\mu^{2}_{E}(\{a,b\},\{a,b,c\})=1/2\hskip 14.22636pt\mu^{2}_{E}(\{a\},\{a,b,c\})=1/4$
    $\mu^{2}_{E}(\{a,c\},\{a,b,c\})=\mu^{2}_{E}(\{a,b,c\},\{a,b,c\})=1/8$
    $\mu^{2}_{E}(\{a,b\},\{a,b\})=1/2\hskip 14.22636pt\mu^{2}_{E}(\{a,c\},\{a,c\})=1/2$
    $\mu^{2}_{E}(\{b\},\{b,c\})=1/8\hskip 14.22636pt\mu^{2}_{E}(\{b,c\},\{b,c\})=1/8\hskip 14.22636pt\mu^{2}_{E}(\{c\},\{b,c\})=1/2$
[…]
    With the above attention function which is also monotonic and the preference $c\succ b\succ a$ we can also rationalize the observed choice probabilities.
Monotonicity in itself is not enough to refine our identification results, however, further restrictions will be. In this direction, I characterize the RAM-UR-IRA by restricting the attention function to be of the Independent Random Attention (IRA) form as in Manzini and Mariotti (2014). The RAM-UR-IRA then allows for a unique identification of the underlying preferences and the attention function.
4 Random Attention Model with Unobserved References: Independent Random Attention
In this section I assume that the attention function in the RAM-UR model is of an independent random type in the sense that there exists an attention parameter $\gamma:X\rightarrow(0,1]$ such that the attention function takes the form: $\mu_{IRA}(D,A)=\prod_{x\in D\subseteq A}\gamma(x)\prod_{y\in A\setminus D}(1-\gamma(y))$ for all $A\in\mathcal{X}\setminus\phi$. Here attention privilege manifests in the form of alternatives with the attention parameter to be one, which is the key difference between this form and the attention function of the IRA form introduced by Manzini and Mariotti (2014) (hereafter referred to as the IRA model). The IRA restricts the attention parameter to be strictly between zero and one while we do not impose such a restriction. This makes the IRA model to be nested in the RAM-UR-IRA.
Here the set of attention-privileged alternatives $E$ is just the set containing all alternatives $x\in X$ such that $\gamma(x)=1$. Since the $\gamma$ function takes strictly positive values for all alternatives, the RAM-UR-IRA is a special case of the RAM-UR as the full support assumption will hold.
Definition 7
A stochastic choice function $p$ is a random attention model with unobservable references and independent random attention (RAM-UR-IRA) if there exists a strict total order $\succ$ over $X$ and an attention parameter $\gamma:X\rightarrow(0,1]$ such that for all $A\in\mathcal{X}\setminus\phi$ and $x\in A$:
$p(x,A)=\gamma(x)\prod_{y\in A;y\succ x}(1-\gamma(y))$
Imposing the above additional structure significantly changes how the probabilities can be defined in the RAM-UR-IRA model. Even though it is a special case of RAM-UR and nests [^11] By nesting the IRA (rational) I mean that a special case of RAM-UR-IRA: the one where $E=\phi$ ($E=X$) will also be an IRA (rational) model. both the IRA as well as classical rationality, it is mathematically a small departure from the IRA model: It allows $\gamma$ to take value one for some alternatives. In what follows, I provide a characterization of this model by using axioms similar to the characterizations of both the IRA and the RAM-UR and discuss the identification results.
4.1 Characterization
I characterize the RAM-UR-IRA using the following five independent axioms:
R-Asymmetry (R-ASYM): For all $A\in\mathcal{X}\setminus\phi$ and $a,b\in A$ with $p(a,A)>0$, $p(b,A)>0$: $\frac{p(a,A\setminus\{b\})}{p(a,A)}\neq 1$ implies $\frac{p(b,A\setminus\{a\})}{p(b,A)}=1$.
This axiom is similar to the i-asymmetry axiom that characterizes the IRA (cf. Manzini and Mariotti (2014)). However, it restricts the denominator to be non-zero. Such a restriction was not needed in the IRA characterization, since all probabilities are assumed to be positive in that model. The axiom says that if removing an alternative $b$ changes the choice probability of another, say $a$ (given the positive denominator constraints), then removing $a$ would not change the choice probability of $b$.
R-Independence (R-IND): (1) For all $A,B\in\mathcal{X}\setminus\phi$ and $a,b\in A\cap B$ with $p(a,A)>0,p(a,B)>0$: $\frac{p(a,A\setminus\{b\})}{p(a,A)}=\frac{p(a,B\setminus\{b\})}{p(a,B)}$.
(2) For all $A,B\in\mathcal{X}\setminus\phi$ and $b\in A\cap B$ with $p(a^{*},A)>0,p(a^{*},B)>0$: $\frac{p(a^{*},A\setminus\{b\})}{p(a^{*},A)}=\frac{p(a^{*},B\setminus\{b\})}{p(a^{*},B)}$.
This axiom is similar to the i-independence axiom in the characterization of the IRA (cf. Manzini and Mariotti (2014)), with the same constraints regarding non-zero probability when needed. It says that the ratio effect of removing an alternative on another alternative’s choice probability (subject to the positive denominator constraints), is menu-independent, i.e. it is the same whatever be the menu from which it is removed.
Non-Triviality (NT): $p(x,\{x\})>0$ for all $x\in X$.
This axiom implies that each alternative has a value on its own in the sense that if this axiom is not satisfied and we have an alternative that is not chosen in a singleton menu with positive probability, then other axioms will ensure that this alternative will not be chosen with a positive probability in any menu. Note that EDA introduced in Section 2.2 implies NT.
Existence of a Dominant Alternative* (EDA*): For all $A\in\mathcal{X}\setminus\phi$: $p(a^{*},A)=0$, implies $p(a^{*},\{x\})=0$ for some $x\in A$.
This axiom is similar in spirit to the EDA axiom used in the characterization of RAM-UR. However, we only need a check for the default alternative as opposed to checking for all alternatives that are chosen with zero probability in a menu. Note again that EDA implies EDA*, which means that NT and EDA* together are weaker than EDA.
Regularity (REG): For all $A,B\in\mathcal{X}\setminus\phi$: $A\subset B$ implies $p(a,A)\geq p(a,B)$ for all $a\in A^{*}$.
This is the standard regularity axiom which states that an alternative’s choice probability weakly reduces in a larger menu. This property is in general not satisfied by the RAM-UR. If we consider the IRA model, the i-asymmetry and the i-independence axioms imply the regularity property. However, that is not the case with the other four axioms utilized to characterize the RAM-UR-IRA. Thus, the property is used as an axiom for the characterization. In the absence of non-triviality, this axiom implies that any trivial alternative [^12] An alternative $x\in X$ here is called trivial if $p(x,\{x\})=0$. will not be chosen with a positive probability in any menu.
Theorem 2
A stochastic choice function is representable by a RAM-UR-IRA model if and only if it satisfies R-ASYM, R-IND, NT, EDA* and REG. Moreover, the representation is unique.
Representation uniqueness means that the set of references, the underlying preferences, and the underlying attention parameters are uniquely identified. This uniqueness result is due to the additional structure imposed on the attention function and to the availability of data in menus with no reference points.
The proof of Theorem 2 essentially partitions the menus into those with no references and ones with at least one reference. Using the axioms above and the Manzini and Mariotti (2014) result on the set of menus with no reference, it identifies the underlying attention parameter and partial preferences among non reference alternatives (precisely the part of the preferences that were unidentified in the RAM-UR). Since the RAM-UR-IRA is a special case of the RAM-UR, the rest of the identification is easily done similar to the general model. The proof shows that the fitting of the identified preferences through the non-reference menus and the preferences identified through the menus with references does indeed give us a strict total order. Finally, using the first two axioms, we obtain the required functional form.
An advantage of a model like the RAM-UR is that it is in the realm of the stochastic choice function, which is why a nestedness analysis with some classical models can be undertaken. Both the RAM-UR and the RAM-UR-IRA nest rational choice as a special case, i.e. when the set of references is the entire set of alternatives. They both also nest the IRA, i.e. when the structure of the attention function is the IRA structure and the set of references is empty. Finally, the RAM-UR-IRA is also representable by a Random Utility Model (RUM) in a similar spirit to how the IRA can be. This is not true for the RAM-UR since it can violate regularity which is necessary for a RUM representation. A brief discussion on the RUM representation of the RAM-UR-IRA is provided in Appendix 2.
5 Literature Overview
The idea that limited attention affects behavior and choice is well accepted and widely documented in the literature. Simon’s (1956) model of satisficing choice is one of the earliest models that can be motivated by limited attention. In this model, the heuristic is that the DM stops at the first alternative which is “good enough”. A DM might utilize such a heuristic because of lack of time or attention. Lleras et al. (2017) model “Choice Overload” which leads to worse outcomes in larger menus, the reason being that good options might fail to grab attention in larger sets due to an overload of available options. Masatlioglu et al. (2012) introduce the concept of the attention filter which incorporates the idea that removing an unconsidered alternative does not change the consideration set in a deterministic choice setting where the DM is attention constrained. Dean et al. (2017) is another paper that incorporates limited attention and also the effect of reference alternatives.
Models of stochastic choice ventured into modelling the variability observed in choices due to limited attention. Manzini and Mariotti (2014) was the seminal paper that filled this void. Since then there has been a substantial body of literature combining both limited attention and stochastic choice. A plethora of papers are closely related to Manzini and Mariotti (2014), for instance, Brady and Rehbeck (2016)’s feasibility model is a generalization of the Manzini and Mariotti (2014), though their motivation is not limited attention, but limited feasibility. Another one is Horan (2019) which looks at the IRA without observing the probability of default. Cattaneo et al. (2020) introduces the general idea of a random attention model. Kovach and Suleymanov (2023) uses the idea of a RAM in the context of reference-dependent choice and models a family of attention functions, each indexed by the reference point, and looked at different special cases like the Constant Random Attention, Luce Random Attention and the Independent Random Attention. They discussed the nestedness properties among these special cases. They were unable to conduct the nestedness analysis with respect to the classical stochastic choice models because in their model, the primitives were menu-reference pairs, which placed it outside the realm of stochastic choice models.
The idea of reference-dependent choice has also been the focus of several papers. Masatlioglu and Ok (2005) and Masatlioglu and Ok (2014) are seminal papers in this area. These authors were one of the first to model status quo bias in choice, which is a widely observed behavior in experiments on individual choice [^13] Masatlioglu and Ok (2005) provides references for such experiments.. Sagi (2006) imposes a no-regret condition on the reference dependent preferences and Tapkı (2007) allows for incomplete reference dependent preferences. Note that all mentioned works model the effect of the status quo on preferences. However, it is possible that the status quo might affect attention as well. Dean et al. (2017) looks at status quo bias and limited attention together, allowing for both reference-dependent preferences and attention. Kovach and Suleymanov (2023), as discussed before, constructs a model where references affect only the attention. Bossert and Sprumont (2009) provides another model where reference points don’t affect preferences. They model a deterministic choice function where the choice is consistent with not choosing a worse alternative than the reference. They study the existence of a linear order such that the choice from each menu-reference pair is at least as good as the reference point, though not necessarily the maximum [^14] The best alternative according to the preference ordering.. The RAM-UR does have this non-deteriorating property since references are attention privileged, so in any menu, an alternative worse than the reference is never chosen.
In the literature discussed above, a general theme is that the primitives include the reference point. It is natural to challenge the assumption of the observability of the reference point. Ok et al. (2015) and Tserenjigmid (2019) are two relevant papers here. However, in Tserenjigmid (2019) the references are not alternatives but a vector of attributes which depend upon the minimum of those attributes among the alternatives in any menu. It is thus, menu dependent as well. Kovach and Suleymanov (2023) is the first paper to study reference dependence in a stochastic setting, which makes the current work the first to study unobservable references in this setup.
We have assumed menu-independence with respect to the references. In Kovach and Suleymanov (2023), the input to the problem is menu-reference pairs so the question of menu independence was moot. However, it is possible to allow for menu-dependent unobservable references as a possible extension to the model in the current paper. This extension obviously has bite, i.e., we cannot explain any stochastic choice function through such a model. For example, any stochastic choice function that allows for alternative $x$ to be chosen with certainty in a menu where $y$ is present and allows alternative $y$ to be chosen with certainty in a menu where $x$ is present cannot be a reference-dependent RAM even with menu-dependent references.
The literature on marketing and consumer research focused on consideration sets and the impact of memory and attention on whether certain brands enter consumers’ consideration sets. Nedungadi (1990) shows that whether an alternative is recalled while forming consideration sets is not necessarily correlated with how highly that alternative is evaluated by the consumer, i.e. attention might be independent of preferences. Advertising often plays a crucial role in increasing an alternative’s possibility of consideration, by increasing its brand recall, or in the terminology of our paper, making it a reference. It is also possible that higher advertising, and thus, more attention, increases the probability of choice of another alternative. This is consistent with both the experiments of Huber et al. (1982) and Nedungadi (1990), and something which can be observed in the model described in this paper. If a reference is added to a menu, the probability of choice of a different alternative can increase. The model described in this paper also seems relevant for more important decision problems like university choice for high school students (Dawes and Brown (2005), Rosen et al. (2008)) and political choice (Wilson (2008)). In these contexts, a fixed set of reference alternatives also makes sense.
Finally, the presence of a default alternative in our model requires some justification. In models of limited attention, the default is used to model the possibility of choice avoidance, delaying choice, which is quite a well-established empirical observation (Dhar (1997)). Horan (2019) argued that it might not be possible to observe the probability of no choice from empirical choice data, and provided a characterization of the IRA by Manzini and Mariotti (2014) without the need for observing such probabilities. However, note that in this model, the default is a de facto reference, and as such, any menu where another reference alternative is present, the default is not relevant. A reference in any menu is the alternative to which any alternative which is considered is always evaluated against. The separation of consideration and evaluation elements of the choice means that the default then becomes the reference for salience. The probability of choosing to abstain is the probability that none of the alternatives will be salient enough to draw attention. If there are other references present, those that are salient enough to always draw attention, the default is redundant.
Appendix
Appendix 1
Proof of Theorem 1
Necessity:
EDA: Note that an alternative $x$ is never chosen from a menu if whenever it is paid attention to, a better alternative is paid attention to as well, which is only possible if a better alternative, say $y$ is attention privileged. In a binary menu including $x$ and $y$, $y$ will again be attention privileged so EDA will hold.
C-WARP: if $y$ is always chosen in a menu, then it is attention-privileged, and better than any other alternative because if it wasn’t, the full support assumption would mean that the other alternative would be chosen with positive probability. So, C-WARP will be satisfied.
EXP: If the antecedent of the axiom holds, then both $y$ and $x$ are attention-privileged, and every other alternative in $A^{*}$ and $B^{*}$ is less preferred to one of the two, so will not be chosen with positive probability. If $x$ is preferred to $y$ then it will always be chosen, and vice versa. So, EXP holds.
Sufficiency:
I take the set $\bar{E}=\{x\in X|p(x,\{x\})=1\}$ to be the revealed set of reference alternatives. If This set is empty EDA ensures that for any $A\in\mathcal{X}\setminus\phi$, $p(a,A)>0$ for all $a\in A^{*}$. In this case, we can construct an attention function with $E=\bar{E}=\phi$ and an arbitrary preference. Note that with an empty set of references, the RAM-UR is just a RAM, so it suffices to show a RAM representation. For this, I take an arbitrary preference and construct an attention function as follows: For any alternative the probability that it is chosen is taken to be the sum of the value that the attention function takes on subsets where that alternative is the best. One way to do it is divide the probability of an arbitrary alternative $x$ in a menu $A$ to all sets in $\{D\subseteq A:x\succ y,\forall y\in D\setminus\{x\}\}$ equally ($\succ$ is the arbitrary preference). Note that since all alternatives (and default) are chosen with positive probability in all menus such a construction will give us a valid attention function. Given the constructed attention function, the observed choices are represented by a RAM with the attention function and the arbitrary preferences. Such a construction will be one possible RAM representation. We can have multiple representations if we change how we assign the probabilities. Thus, we have a representation when $\bar{E}$ is empty. For the rest of the proof I assume that $\bar{E}$ is non-empty.
In such a case, I define the binary relation $P$ over $X$ as follows:
  1. 1.
    For $x\in\bar{E}$ and $y\in X^{*}$ with $x\neq y$: $xPy$ iff $p(x,\{x,y\})=1$.
  2. 2.
    For $x\notin\bar{E}$ and $y\in\bar{E}$: $xPy$ iff $p(x,\{x,y\})>0$.
  3. 3.
    For $x,y\notin\bar{E}$: $xPy$ iff there exists a $z\in\bar{E}$ such that $xPz$ and $zPy$.
Lemma 1
$P$ is a strict partial order over $X$.
Proof. To show the asymmetry of $P$, we need to show that it is antisymmetric and irreflexive. I take it case-wise:
If $x\in\bar{E}$, then I haven’t defined $xPx$. If $x\neq y$, then $xPy$ implies $p(x,\{x,y\})=1$. So, $yPx$ is not possible since $p(y,\{x,y\})=0$, and $yPx$ cannot be the case. So for this case, there is no violation of irreflexivity and antisymmetricity.
If $x\notin\bar{E}$, then $xPx$ is only possible if there exists a $z\in\bar{E}$ such that $xPz$ and $zPx$ but that is not possible as above, so $P$ is irreflexive. Now, if $y\neq x$ such that $y\in\bar{E}$, then $xPy$ implies $p(x,\{x,y\})>0$, and thus, $yPx$ is not possible. Finally, if $y\neq x$ such that $y\notin\bar{E}$, then $xPy$ implies the existence of $z\in\bar{E}$ such that $p(x,\{x,z\})>0$ and $p(z,\{z,y\})=1$. If it were possible that there exists $w\in\bar{E}$ such that $p(w,\{w,x\})=1$ and $p(y,\{w,y\})>0$, then by EXP we have either $p(w,\{w,x,y,z\})=1$ or $p(z,\{w,x,y,z\})=1$ in contradiction to C-WARP. So, $P$ is asymmetric.
Now I show the transitivity of $P$. I will again argue casewise, in each of the cases having $xPy$ and $yPz$ for $x,y,z\in X$:
  • Case 1: $x\in\bar{E}$, $y\in\bar{E}$ and $z\in X^{*}$.
    By the definition of $P$ we have $p(x,\{x,y\})=1$ and $p(y,\{y,z\})=1$. EXP and C-WARP give us $p(x,\{x,y,z\})=1$, applying C-WARP again gives us $p(x,\{x,z\})=1$, so we have $xPz$.
  • Case 2: $x\in\bar{E}$, $y\notin\bar{E}$ and $z\in\bar{E}$.
    We have $p(x,\{x,y\})=1$ and $p(y,\{y,z\})>0$. Note that $p(z,\{z\})=1$ since $z\in\bar{E}$. So, by EXP, either $p(x,\{x,y,z\})=1$ or $p(z,\{x,y,z\})=1$, but the latter along with $p(y,\{y,z\})>0$ violates C-WARP, so we must have $p(x,\{x,y,z\})=1$ which by C-WARP gives us $p(x,\{x,z\})=1$ and thus, $xPz$.
  • Case 3: $x\in\bar{E}$, $y\notin\bar{E}$ and $z\notin\bar{E}$.
    We have $p(x,\{x,y\})=1$, as well as the existence of $w\in\bar{E}$ such that $yPw$ and $wPz$. Note that $w\neq x$ since $P$ is asymmetric. Looking at the menu $\{x,w,z\}$ we have $p(x,\{x,w,z\})=1$ because by Case 2 we know that $xPw$ and by EXP we know that either $p(x,\{x,w,z\})=1$ or $p(w,\{x,w,z\})=1$ holds. By C-WARP, $xPz$ follows.
  • Case 4: $x\notin\bar{E}$, $y\in\bar{E}$ and $z\in\bar{E}$.
    We have $p(x,\{x,y\})>0$ and $p(y,\{y,z\})=1$. If $p(x,\{x,z\})=0$, then we have $p(z,\{x,z\})=1$ by EDA. By EXP, either $p(y,\{x,y,z\})=1$ or $p(z,\{x,y,z\})=1$ must be the case, so $p(x,\{x,y,z\})=0$. By C-WARP, it must be that $p(z,\{x,y,z\})=1$, and applying C-WARP again we get $p(z,\{y,z\})=1$ in contradiction to $yPz$. So, it must be that $p(x,\{x,z\})>0$ and thus, $xPz$ follows.
  • Case 5: $x\notin\bar{E}$, $y\in\bar{E}$ and $z\notin\bar{E}$.
    We have $xPy$, $yPz$ and $y\in\bar{E}$, so by definition $xPz$ follows.
  • Case 6: $x\notin\bar{E}$, $y\notin\bar{E}$ and $z\in\bar{E}$.
    We have $xPy$ and thus, there exists $w\in\bar{E}$ such that $xPw$ and $wPy$. By Case 2 we have $wPz$, and by Case 4 we have $xPz$.
  • Case 7: $x\notin\bar{E}$, $y\notin\bar{E}$ and $z\notin\bar{E}$.
    We have $xPy$ and thus, there exists $w\in\bar{E}$ such that $xPw$ and $wPy$. So, by Case 3 we have $wPz$. By Case 5 $xPz$ follows.
In other words, the binary relation $P$ defined above is transitive. Note that P is complete when it comes to comparisons between an alternative in $\bar{E}$ and any other alternative. The only comparisons it could be silent on are between alternatives not in $\bar{E}$.
Given that $P$ is a strict partial order we can extend it to a strict total order [^15] A strict total order is an asymmetric, transitive and complete binary relation over $X$. (Szpilrajn (1930)). Note that such an extension need not be unique. Fix $\succ^{*}$ to be such an extension, that is, for all $x,y\in X$ we have that:
$xPy\implies x\succ^{*}y$
Next, I define the set $D_{x}$ to be the largest subset of $X$ such that $x$ is the $\succ^{*}$-maximal element.
Now I will specify the revealed attention function $\mu^{*}_{\bar{E}}:\mathcal{X}\times\mathcal{X}$ such that:
$p(x,A)=\sum_{D\in\{D:[x\cup(A\cap\bar{E})]\subseteq D\subseteq D_{x}\cap A\}}\mu^{*}_{\bar{E}}(D,A)$
We can have multiple attention functions that satisfy this requirement, I will show one way we can construct a revealed attention function.
Let $y$ be the $\succ^{*}$-maximal element in $A\cap\bar{E}$. Then any alternative $x\in A$ such that $y\succ^{*}x$ will have zero probability because of C-WARP. The set of possible sets $D$ that satisfy the requirement, $\{D:[x\cup(A\cap\bar{E})]\subseteq D\subseteq D_{x}\cap A\}$ is empty, so any attention function $\mu^{*}_{\bar{E}}$ will satisfy the above equality.
For an arbitrary element $x\in A$ which is $\succ^{*}$-better than $y$, EDA and C-WARP say that they will have positive probability in the menu. Similar to the construction by Kovach and Suleymanov (2023) we can ensure the existence of one such attention function by the following procedure: Let $n(x,A):=|\{D:[x\cup(A\cap\bar{E})]\subseteq D\subseteq D_{x}\cap A\}|$. Then we can specify $\mu^{*}_{\bar{E}}(D,A)=\frac{p(x,A)}{n(x,A)}$ for all $D\in\{D:[x\cup(A\cap\bar{E})]\subseteq D\subseteq D_{x}\cap A\}$. This gives us the required functional form. The set $\bar{E}$ is the revealed set of reference points which is unique. $\succ^{*}$ is the identified preferences unique up to the strict partial order $P$. $P$ uniquely gives us the preference of each alternative with respect to each reference points and the transitive closures that it implies.
Proof of Theorem 2
Necessity:
R-ASYM: If there exist $A\in\mathcal{X}\setminus\phi$ and $x,y\in A$ with $p(x,A)>0$ and $p(y,A)>0$, then either none of $x$ and $y$ are references, in which case the axiom is true because of the functional form, as proved in Manzini and Mariotti (2014), or exactly one of them is a reference, and the other is preferred to it. In the latter case, the antecedent is true for the reference, since the probability of choice of the less preferred alternative (the reference) will increase if the other is removed, but removing the reference will not affect the choice probability of the better alternative.
R-IND: If there exists $A,B\in\mathcal{X}\setminus\phi$, $a,b\in A\cap B$ with $p(a,A)>0$ and $p(a,B)>0$, the functional form is such that the R-IND will hold. Same is true if instead of $a$ we consider the default $a^{*}$. The proof is identical to the proof of the necessity of i-independence in Manzini and Mariotti (2014).
NT: This axiom is trivially satisfied given the functional form of the RAM-UR-IRA.
EDA*: If $p(a^{*},A)=0$, for some $A\in\mathcal{X}\setminus\phi$ then there exists at least one alternative $x\in A$ such that $\gamma(x)=1$. Then $p(a^{*},\{x\})=0$ follows.
REG: Consider arbitrary $A\subseteq B\subseteq X$, then the functional form for any $x\in A$ gives us that $p(x,A)=\gamma(x)\prod_{y\in A;y\succ x}(1-\gamma(y))$. Similarly we get $p(x,B)=\gamma(x)\prod_{y\in B;y\succ x}(1-\gamma(y))$ for the set $B$. We can rewrite this as: $p(x,B)=\gamma(x)\prod_{y\in A;y\succ x}(1-\gamma(y))\prod_{y\in B\setminus A;y\succ x}(1-\gamma(y))$. Since all the attention parameters lie between zero and one, it is clear that $p(x,B)\leq p(x,A)$. So, REG is fulfilled.
Sufficiency:
I take the set $\bar{E}=\{x\in X:p(x,A)=1\hskip 1.42262pt\text{for some}\hskip 1.42262ptA\in\mathcal{X}\setminus\phi\}$ to be the revealed set of reference alternatives.
Again, this set could be empty. If this set is empty, then EDA* and NT ensure that $p(a,A)>0$ for all $A\in\mathcal{X}\setminus\phi$ and $a\in A^{*}$. Now the R-ASYM and the R-IND will be equivalent to the i-ASYM and i-IND by Manzini and Mariotti (2014), which gives us the required equivalence with the IRA model [^16] An IRA model is a RAM-UR-IRA model with $E=\phi$.. For the remainder of the proof, I assume non-emptyness of $\bar{E}$.
I will use the notation IRA to denote any reference to Manzini and Mariotti (2014) in this proof. We define two more properties RIIA and EXP, which the above axioms together imply:
Reference-Independence of Irrelevant Alternatives (RIIA): For all $A\in\mathcal{X}\setminus\phi$ and $a\in A$: $p(a,A)=0$ implies $p(b,A\setminus\{a\})=p(b,A)$ for all $b\in A\setminus\{a\}$.
Claim 1: If a stochastic choice function $p:X\times\mathcal{X}\setminus\phi\rightarrow[0,1]$ satisfies REG, then it satisfies RIIA.
Proof: By REG, for any alternative $c\in A^{*}\setminus\{a\}$ such that $p(c,A)>0$, we have $p(c,A)\leq p(c,A\setminus\{a\})$ and since summing over $p(c,A)$ for all $c\in A^{*}$ gives us one, they must be equal.
Expansion (EXP): For all $A,B\in\mathcal{X}\setminus\phi$, $x_{1}\in A,x_{2}\in B$: $p(x_{1},A)=p(x_{2},B)=1$ implies either $p(x_{1},A\cup B)=1$ or $p(x_{2},A\cup B)=1$.
Claim 2: If a stochastic choice function $p:X\times\mathcal{X}\setminus\phi\rightarrow[0,1]$ satisfies R-ASYM, R-IND and REG, then it satisfies EXP.
Proof: By REG, $p(y,A\cup B)=0$ for all $y\in A\cup B)^{*}\setminus\{x_{1},x_{2}\}$. Suppose $p(x_{1},A\cup B)>0$ and $p(x_{2},A\cup B)>0$. Then by R-IND, we have $\frac{p(x_{1},A\cup B\setminus\{x_{2}\})}{p(x_{1},A\cup B)}=\frac{p(x_{1},\{x_{1}\})}{p(x_{1},\{x_{1},x_{2}\})}$ and similarly for $x_{2}$. But $p(x_{1},\{x_{1}\})=p(x_{2},\{x_{2}\})=1$ by REG. So, $\frac{p(x_{1},A\cup B\setminus\{x_{2}\})}{p(x_{1},A\cup B)}>1$ and $\frac{p(x_{2},A\cup B\setminus\{x_{1}\})}{p(x_{2},A\cup B)}>1$ which contradicts R-ASYM.
I proceed with the proof of Theorem 2 and start with the following two lemmas.
Lemma 2
$A\cap\bar{E}=\phi$ for some $A\in\mathcal{X}\setminus\phi$ implies $p(a,A)>0$ for all $a\in A$.
Proof. Suppose not. Then there exist a menu $A\in\mathcal{X}$ with $A\cap\bar{E}=\phi$ and $a\in A$ such that $p(a,A)=0$.
By NT and RIIA there exists $x\in A$ such that $p(x,A)>0$.
By RIIA, $\frac{p(x,A\setminus\{a\})}{p(x,A)}=1$.
Repeatedly applying RIIA we can remove irrelevant alternatives [^17] An alternative $a\in A\in\mathcal{X}\setminus\phi$ is irrelevant with respect to $A$ if $p(a,A)=0$. from A to reach the menu $A^{{}^{\prime}}=\{x:x\in A,p(x,A)>0\}$. Note that $p(x,A)=p(x,A^{{}^{\prime}})$ for all $x\in A^{{}^{\prime}}$.
Also note that since $A^{\prime}\subseteq A$, $A^{\prime}\cap\bar{E}=\phi$. EDA* implies then $p(a^{*},A^{\prime})>0$. Due to the summation of choice probabilities, we have $p(a^{*},A^{\prime})=p(a^{*},A)$. Since $a\in A$ is such that $p(a,A)=0$ we have $p(a^{*},A)=p(a^{*},A\setminus a)$
Using R-IND and NT, we have $11=\frac{p(a^{*},A\setminus\{a\})}{p(a^{*},A)}=\frac{p(a^{*},\phi)}{p(a^{*},a)}>$ 1
which is a contradiction.
Denote by $\mathcal{X}^{{}^{\prime}}$ the set of menus which contain only the elements not in $\bar{E}$. I will first look only at the menus in the subsystem $\mathcal{X}^{{}^{\prime}}$. Lemma 2 indicates that for any menu $A\in\mathcal{X}^{{}^{\prime}}$, $p(a,A)>0$ for all $a\in A^{*}$.
Moreover, R-ASYM and R-IND reduce down to the i-asymmetry and the i-independence condition in the characterization of IRA, so in the subsystem $\mathcal{X}^{{}^{\prime}}$, the required functional form equivalence is achieved due to the characterization result by Manzini and Mariotti (2014). Also, the preferences are elicited using the i-regularity property which is also satisfied over this subsystem. Let the preferences elicited in such a way over alternatives in $X^{{}^{\prime}}$ be $\succ^{{}^{\prime}}$. Obviously, over this set the preference is a strict linear order.
Lemma 3
$A\cap\bar{E}\neq\phi$ for some $A\in\mathcal{X}\setminus\phi$ implies $p(a^{*},A)=0$.
Proof. Suppose not. Then there exists $A\in\mathcal{X}\setminus\phi$ such that $A\cap\bar{E}\neq\phi$ and $p(a^{*},A)>0$. For $x\in A\cap\bar{E}$, we have by REG that $p(a^{*},\{x\})>0$ or $p(x,\{x\})\neq 1$. Since $x\in\bar{E}$, $p(x,B)=1$ for some $B\in\mathcal{X}\setminus\phi$. Applying REG leads to $p(x,\{x\})=1$ which is a contradiction.
I will now construct a binary relation over alternatives in $X$ and combine it with the preference $\succ^{{}^{\prime}}$ generated over $X^{{}^{\prime}}$ to form a preference over all alternatives. I will then show that this combination is a strict total order over all alternatives in $X$. For this, let us define the binary relation $\succ^{{}^{\prime\prime}}$ over $X$ as follows:
For all $a\in\bar{E}$ and $b\in X$: $p(a,\{a,b\})=1$ implies $a\succ^{{}^{\prime\prime}}b$, and $p(a,\{a,b\})\neq 1$ implies $b\succ^{{}^{\prime\prime}}a$.
I can put these two binary relations together to get another binary relation: $\succ=\succ^{{}^{\prime}}\cup\succ^{{}^{\prime\prime}}$.
Claim 3: The binary relation $\succ$ is a strict total order over $X$.
Proof. Obviously, $\succ$ is complete by definition. Among the alternatives in $X^{{}^{\prime}}$ it is clearly asymmetrical as well. We need to show asymmetry among alternatives where at least one of the alternatives is in $\bar{E}$.
Consider an arbitrary $x\in\bar{E}$ and arbitrary $y\in\bar{E}$ preferences are only defined through $\succ^{{}^{\prime\prime}}$ and is asymmetric by definition.
Suppose $a\in\bar{E}$ and $b\in\bar{E}$. By EXP we have either $p(a,\{a,b\})=1$ or $p(b,\{a,b\})=1$. Without loss of generality, if $p(a,\{a,b\})=1$, then $a\succ b$ but $b\nsucc a$. So, $\succ$ is asymmetric.
Now we show that $\succ$ is transitive as well. For this let us look at the following cases.
Case 1: $a\notin\bar{E}$, $b\notin\bar{E}$, $c\notin\bar{E}$, $a\succ b$ and $b\succ c$.
The binary relation $\succ$ here is defined entirely by $\succ^{{}^{\prime}}$ which is transitive as shown in the characterization of IRA, so transitivity of $\succ$ follows.
Case 2: $a\in\bar{E}$, $b\in\bar{E}$, $c\in X$, $a\succ b$ and $b\succ c$.
We have $p(a,\{a,b\})=1$ and $p(b,\{b,c\})=1$. By EXP, either $p(a,\{a,b,c\})=1$ or $p(b,\{a,b,c\})=1$ holds.
Suppose, $p(b,\{a,b,c\})=1$. Then by REG, $p(b,\{a,b\})=1$ follows. However, this contradicts the fact that $a\succ b$ and thus, $p(a,\{a,b,c\})=1$ should hold.
By REG again, we have $p(a,\{a,c\})=1$ and thus, $a\succ c$.
Case 3: $a\in\bar{E}$, $b\notin\bar{E}$ and $c\in\bar{E}$, $a\succ b$ and $b\succ c$.
We have $p(a,\{a,b\})=1$ and $p(b,\{b,c\})>0$ (otherwise, $p(a^{*},\{b,c\})=1$ in contradiction to Lemma 3).
Since $c\in\bar{E}$, we have by REG that $p(c,\{c\})=1$ holds. Using EXP, we have either $p(a,\{a,b,c\})=1$ or $p(c,\{a,b,c\})=1$.
If $p(c,\{a,b,c\})=1$, then by REG we have $p(c,\{b,c\})=1$ which is not true by assumption. So, $p(a,\{a,b,c\})=1$ should be the case.
By REG, we have $p(a,\{a,c\})=1$ and thus, $a\succ c$ follows.
Case 4: $a\in\bar{E}$, $b\notin\bar{E}$ and $c\notin\bar{E}$, $a\succ b$ and $b\succ c$.
We have $p(a,\{a,b\})=1$ and $\frac{p(b,\{b\})}{p(b,\{b,c\})}=1$ (which holds due to $i-asymmetry^{*}$ [^18] This condition is defined by Manzini and Mariotti (2014) which says that the implication in i-asymmetry goes both ways. They show that it is implied by their axioms. which holds over the menus $\{b\}$ and $\{b,c\}$ since they are part of the subsystem $\mathcal{X}^{{}^{\prime}}$ where all the IRA results apply).
Claim: $p(b,\{a,b,c\})=0$.
Proof. If not, then $p(b,\{a,b,c\})>0$. By REG, $p(b,\{a,b\})>0$ which is a contradiction.
By RIIA and the above claim $p(c,\{a,c\})=p(c,\{a,b,c\})$. Suppose $p(c,\{a,b,c\})>0$, then by R-IND, $1=\frac{p(c,\{a,c\})}{p(c,\{a,b,c\})}=\frac{p(c,\{c\})}{p(c,\{b,c\})}$ but $\frac{p(c,\{c\})}{p(c,\{b,c\})}\neq 1$ because of R-ASYM. Thus, it must be that $p(c,\{a,b,c\})=0$.
By RIIA, we have $p(c,\{a,c\})=p(c,\{a,b,c\})=0$. Applying Lemma 3, we get $p(a,\{a,c\})=1$. Thus, $a\succ c$ holds.
Case 5: $a\notin\bar{E}$, $b\notin\bar{E}$ and $c\in\bar{E}$, $a\succ b$ and $b\succ c$.
We have $\frac{p(b,\{b\})}{p(b,\{a,b\})}>1$ (by i-regularity of the IRA result on the subsystem $\mathcal{X}^{{}^{\prime}}$) and $p(b,\{b,c\})>0$ (by Lemma 3).
Claim: $p(a,\{a,b,c\})>0$.
Proof. Suppose not. Then $p(a,\{a,b,c\})=0$. By RIIA, $p(b,\{b,c\})=p(b,\{a,b,c\})>0$. By R-IND, $11=\frac{p(b,\{b,c\})}{p(b,\{a,b,c\})}=\frac{p(b,\{b\})}{p(b,\{a,b\})}>$ 1 which is not possible. So, $p(a,\{a,b,c\})>0$ follows.
By R-IND, we have $\frac{p(a,\{a,c\})}{p(a,\{a,b,c\})}=\frac{p(a,\{a\})}{p(a,\{a,b\})}$. We can use R-IND since by Lemma 1 we have $p(a,\{a,b\})>0$ as well. Note that by NT, we have $p(a,\{a\})>0$ and thus, we must have $p(a,\{a,c\})>0$. Hence, $a\succ c$ follows.
Case 6: $a\notin\bar{E}$, $b\in\bar{E}$ and $c\in\bar{E}$, $a\succ b$ and $b\succ c$.
We have $p(a,\{a,b\})>0$ (by Lemma 3) and $p(b,\{b,c\})=1$.
Claim: $p(a,\{a,b,c\})>0$.
Proof. Suppose not. then $p(a,\{a,b,c\})=0$. By RIIA we have $p(c,\{a,b,c\})=p(c,\{b,c\})=0$ and $p(b,\{a,b,c\})=p(b,\{b,c\})=1$. Applying RIIA again gives us $p(b,\{a,b\})=p(b,\{a,b,c\})=1$ which is not possible, so $p(a,\{a,b,c\})>0$.
By R-IND, $\frac{p(a,\{a,c\})}{p(a,\{a,b,c\})}=\frac{p(a,\{a\})}{p(a,\{a,b\})}$. This implies $p(a,\{a,c\})>0$ because $p(a,\{a\})>0$ and $p(a,\{a,b\})>0$. Thus, $a\succ c$ holds.
Case 7: $a\notin\bar{E}$, $b\in\bar{E}$ and $c\notin\bar{E}$, $a\succ b$ and $b\succ c$.
We have $p(a,\{a,b\})>0$ (by Lemma 3) and $p(b,\{b,c\})=1$.
Claim: $p(a,\{a,b,c\})>0$.
Proof. Suppose not. then $p(a,\{a,b,c\})=0$. By RIIA we have $p(c,\{a,b,c\})=p(c,\{b,c\})=0$ and $p(b,\{a,b,c\})=p(b,\{b,c\})=1$. Applying RIIA again gives us $p(b,\{a,b\})=p(b,\{a,b,c\})=1$ which is not possible, so $p(a,\{a,b,c\})>0$.
Note that $p(c,\{a,b,c\})=0$ by REG since $p(c,\{b,c\}=0$ by assumption.
Since $a\notin\bar{E}$, $p(a,\{a,b,c\})<1$. Since $p(c,\{a,b,c\})=0$ and $p(a^{*},\{a,b,c\})=0$ (by Lemma 3), we have $p(b,\{a,b,c\})>0$.
It follows from $\frac{p(b,\{b,c\})}{p(b,\{a,b,c\}}>1$ and R-ASYM that $\frac{p(a,\{a,c\})}{p(a,\{a,b,c\})}=1$ holds. By R-IND, we have $\frac{p(a,\{a\})}{p(a,\{a,c\})}=\frac{p(a,\{a,c\})}{p(a,\{a,b,c\})}=1$. In the subsystem containing the menu $\{a,c\}$, $i-asymmetry^{*}$ holds and thus, $\frac{p(c,\{c\})}{p(c,\{a,c\})}>1$ leads to $a\succ c$.
In view of the above cases we have that $\succ$ is transitive and thus, a strict total order over $X$.
We now define $\gamma:X\rightarrow(0,1]$ by $\gamma(a)=p(a,\{a\})$ for each $a\in X$. Note that $\gamma(a)=1$ for all $a\in\bar{E}$ and $\gamma(b)\in(0,1)$ for all $b\notin\bar{E}$. $\bar{E}$ will be identified here as the set $E$ of the RAM-UR-IRA.
Claim 4: For all $A\in\mathcal{X}\setminus\phi$ and $a,b\in A$: $p(a,A)>0$, $p(b,\{b\})\neq 1$ and $p(a,A\setminus\{b\})>p(a,A)$ imply $\frac{p(a,A\setminus\{b\})}{p(a,A)}=\frac{1}{1-p(b,\{b\})}$ for all $a\in A^{*}$.
Proof. This claim is similar to the one in the proof of the first theorem in Manzini and Mariotti (2014). Proving this claim will allow us to recover the functional form of the IRA structure.
For $a^{*}$, the required equality translates to $\frac{p(a^{*},A\setminus\{b\})}{p(a^{*},A)}=\frac{1}{1-p(b,\{b\})}$. Note that $\frac{1}{1-p(b,\{b\})}=\frac{p(a^{*},\phi)}{p(a^{*},\{b\})}$. which is true by the implementation of R-IND.
For $a,b\in A$ such that $\frac{p(a,A\setminus\{b\})}{p(a,A)}>1$. Here, RIIA implies that $p(b,A)>0$.
We can apply R-ASYM, to get $p(b,A\setminus\{a\})=p(b,A)$
By REG we get $p(a,\{a,b\})>0$.
By R-IND we have,
$\frac{p(a,A\setminus\{b\})}{p(a,A)}=\frac{p(a,\{a\})}{p(a,\{a,b\})}=\frac{1-p(a^{*},\{a\})}{1-p(b,\{a,b\})-p(a^{*},\{a,b\})}$
. Applying R-ASYM,
$\frac{p(a,A\setminus\{b\})}{p(a,A)}=\frac{1-p(a^{*},\{a\})}{1-p(b,\{b\})-p(a^{*},\{a,b\})}$
If $p(a,\{a\})=1$, then we get the required equation by REG. If not we have $a\notin\bar{E}$ so by EDA* we have $p(a^{*},\{a,b\})>0$. We can then use the following equation from the first part of this proof:
$p(a^{*},\{a\})=\frac{p(a^{*},\{a,b\})}{1-p(b,\{b\})}$
Putting this above we will get the required equality:
$\frac{p(a,A\setminus\{b\})}{p(a,A)}=\frac{1}{1-p(b,\{b\})}$
Now I will look at different set of menus and alternatives to show that we get the required functional form.
For $a\in\bar{E}$, define $A=\{a\}\cup\{x:x\in X,a\succ X\}$. By definition of $\succ$, $p(a,\{a,b\})=1$ for all $b\in A$. So, by REG, $p(a,A)=1=\delta(a)$. As required.
Let $A\in\mathcal{X}$ be a menu such that there exists atleast one $x\in A\cap\bar{E}$. Let $a\in A\cap\bar{E}$ such that $a\succ b$ for any $b\in A\cap\bar{E}\setminus\{a\}$, essentially, $a$ is the most preferred alternative in $A\cap\bar{E}$.
For any alternative $b\in A$ such that $a\succ b$, $p(b,A)=0$ by REG. Also by REG, $p(a^{*},A)=0$ as required.
Claim: In the above case, $p(a,A)>0$.
Proof. Suppose not. Then $p(a,A)=0$.
Since we have assumed $p(a,A)=0$, $p(c,A)=p(c,A\setminus\{a\})$ follows for all $c\in A\setminus\{a\}$. We can iteratively remove all alternatives in $A$ that are worse than $a$ and $a$ itself to reach $A^{{}^{\prime}}=\{b\in A:b\succ a\}$.
Since, $a$ is the most preferred alternative in $\bar{E}$ that is in $A$, $A^{{}^{\prime}}\cap\bar{E}=\phi.$. So by EDA*, $p(a^{*},A^{{}^{\prime}})>0$.
Since $p(b,A)=p(b,A^{{}^{\prime}})$ for all $b\in A^{{}^{\prime}}$ and $p(c,A)=0$ for all $c\in A\setminus(A^{{}^{\prime}})$, $p(a^{*},A)>0$ follows in contradiction to REG.
Claim: If $c\succ a$ and $c\in A$, then $p(c,A)>0$.
Proof. Suppose, there exists $c\in A$ such that $p(c,A)=0$ and $c\succ a$. Since $p(a,A)>0$, we have by R-IND $1<\frac{p(a,\{a\})}{p(a,\{a,c\})}=\frac{p(a,A\setminus\{c\})}{p(a,A)}$ which is incompatible with RIIA. So, it must be that $p(c,A)>0$ for all $c\succ a$.
I can now iteratively apply Claim 4 similar to how it was done in Manzini and Mariotti (2014) to get the required equivalence with the functional form. We have already considered menus without any alternatives in $\bar{E}$ through the subsystem $\mathcal{X}^{{}^{\prime}}$. The rest of the cases have been considered above. $E$ is identified uniquely as the $\bar{E}$ and preferences as specified above are identified uniquely by $\succ$.
Appendix 2
In this section, I discuss the Random Utility Model representation of the RAM-UR-IRA and analyze how the representation changes as the set of reference alternatives $E$ changes.
Denote by $\mathcal{L}(X^{*})$ the set of all strict total orders over $X^{*}$ with typical element $R$. A probability distribution over $\mathcal{L}(X^{*})$ is denoted by $\nu$.
Definition 8
A stochastic choice function $p:X^{*}\times\mathcal{X}\setminus\phi$ has a Random Utility Model (RUM) representation if there exists a distribution $\nu$ over the set of all strict total orders on $X^{*}$, $\mathcal{L}(X^{*})$ such that:
$p(x,A)=\sum_{R\in\mathcal{L}(X^{*}):xRy\forall y\in A^{*}\setminus\{x\}}\nu(R)$
for all $x\in A^{*}$.
The representation is very similar in construction to the one provided by Manzini and Mariotti (2014) for their IRA model which is a RUM too. The construction for the IRA goes as follows:
  • For any alternative $a\in X$, set $\nu(\{R:aRa^{*}\})=\gamma(a)$.
  • If $a\succ b$, then for an $R$ such that $bRa$, $aRa^{*}$ and $bRa^{*}$, set $\nu(R)=0$.
  • For any two alternatives $a$ and $b$, $\nu(\{R:aRa^{*}\hskip 2.84526pt\text{and}\hskip 2.84526ptbRa^{*}\})=\gamma(a)\gamma(b)$.
The above restrictions continue to hold for the RAM-UR-IRA but we will also need more restrictions for our model:
  1. 1.
    A reference is always considered if available. So, we have $\nu(\{R:a^{*}Ra\})=0$ for all $a\in E$.
  2. 2.
    If $a\in E$ and $a\succ b$ for $b\in X$, set $\nu(\{R:b\succ a\})=0$.
(1) here is implied by the first point in the construction by Manzini and Mariotti (2014). It is reiterated to emphasize that all strict orders where the default beats a reference alternative are essentially removed from the support of the RUM.
Note that as we add more alternatives to the set $E$, more and more strict total orders are removed from the support. As $E$ becomes the set $X$, only the strict linear order, which is the preference of the agent is in the support, and the agent is rational. On the other hand, if $E$ is the empty set, the first set of restrictions will be the only ones that apply, and we are back to the Manzini and Mariotti (2014) model. Looking at the RUM representation makes the nestedness properties of the RAM-UR-IRA clearer.
Note finally that the RAM-UR might not be a RUM because it fails to satisfy regularity which is a necessary condition for a stochastic choice function to be a RUM, (Falmagne (1978)). This is the case because of the lack of structure on the attention function due to which for an alternative that is undominated by any reference in a smaller and a bigger menu, we can construct attention probabilities of the subsets where it is the best alternative to be more in the bigger menu than in the smaller menu. This will lead to a RAM-UR model that violate regularity. The lack of structure of the general model allows us to construct attention functions that gives us the means to explain different behavioral observations. One such empirically observed behavioral pattern that it allows us to explain is that the status quo (references) is more likely to be chosen in larger sets (Choice Overload) similar to how Kovach and Suleymanov (2023) showed in their model using the Luce Random Attention structure.